### v2

# Determination of the density in the linear elastic wave equation

Abstract.
We study the inverse boundary value problem for the linear elastic wave equation in three-dimensional isotropic medium. We show that both the Lamé parameters and the density can be uniquely recovered from the boundary measurements under the strictly convex foliation condition.
Key words and phrases:
elastic waves, inverse boundary value problem, density, transverse ray transform
J. Zhai is supported by National Key Research and Development Programs of China (No. 2023YFA1009103), NSFC(No. 12471396), Science and Technology Commission of Shanghai Municipality (23JC1400501)
1. Introduction
The wave equation for the displacement vector $u=(u_{1},u_{2},u_{3})$ in a three-dimensional isotropic elastic body $\Omega$ reads
(1) $(Pu)_{i}=\rho\partial^{2}_{t}u_{i}-\sum_{jk\ell}\partial_{j}(c_{ijk\ell}\partial_{\ell}u_{k})=0,$
where $c_{ijk\ell}=\lambda\delta_{ij}\delta_{k\ell}+\mu(\delta_{ik}\delta_{j\ell}+\delta_{i\ell}\delta_{jk})$. Here we assume that the Lamé parameters $\lambda,\mu$ and the density $\rho$ all belong to $C^{\infty}(\overline{\Omega})$ and
(2) $\mu>0,\quad 3\lambda+2\mu>0,\quad\rho>0,\quad\text{on }\overline{\Omega}.$
We will consider the initial boundary value problem
(3) $\begin{cases}Pu=0,\quad&\text{in }(0,T)\times\Omega,\\ u=f,\quad&\text{on }(0,T)\times\partial\Omega,\\ u=\partial_{t}u=0,\quad&\text{on }\{t=0\}\times\Omega,\end{cases}$
where $\Omega\subset\mathbb{R}^{3}$ is an open domain, and the boundary $\partial\Omega$ is smooth. We will consider the inverse boundary value problem of recovering $\lambda,\mu,\rho$ from the Dirichlet-to-Neumann map
$\Lambda_{\lambda,\mu,\rho}:f\mapsto\sum_{jk\ell}\nu_{j}c_{ijk\ell}\partial_{\ell}u_{k}|_{(0,T)\times\partial\Omega},$
where $\nu$ is the unit outer normal to $\partial\Omega$. The main application of this problem is seismic inversion [1, 22], that is to recover the subsurface geologic structure from seismic data.
The principal symbol of $P$, as a pseudodifferential operator, is
$p_{ik}(t,x,\tau,\xi)=\rho(x)\tau^{2}\delta_{ik}-c_{ijk\ell}(x)\xi_{j}\xi_{\ell}.$
The characteristic set $\{(t,x,\tau,\xi)|\det p(t,x,\tau,\xi)=0\}$ of $P$ consists of two separate sets $\Sigma_{P}=\{(t,x,\tau,\xi)|\tau^{2}-c_{P}^{2}(x)|\xi|^{2}=0\}$ and $\Sigma_{S}=\{(t,x,\tau,\xi)|\tau^{2}-c_{S}^{2}(x)|\xi|^{2}=0\}$, where
$c_{P}=\sqrt{\frac{\lambda+2\mu}{\rho}},\quad\quad c_{S}=\sqrt{\frac{\mu}{\rho}}$
are the wavespeeds for $P$- and $S$- waves. Clearly (2) implies $c_{P}>c_{S}$. The $P/S$-wave paths are then modeled by integral curves (zero-bicharacteristics) $(t(s),x(s),\tau(s),\xi(s))$ of the Hamiltonian vector field associated with the Hamiltonian $\tau^{2}-c^{2}_{P/S}(x)|\xi|^{2}$. The bicharacteristics $(x(s),\xi(s))$ is just the lift-ups of geodesics with respect to the Riemannian metric $g_{P/S}:=c_{P/S}^{-2}g_{E}$ to the phase space, where $g_{E}$ is the Euclidean metric.
The corresponding problem for the scalar wave equation $(\partial^{2}_{t}-\Delta_{g})u=0$ is well studied, where $g$ is a Riemannian metric. It is well known that one can recover $g$ up to an isometry fixing $\partial\Omega$ using the boundary control method [5]. For a different approach, under certain geometric conditions, the problem can also be reduced to some geometric inverse problems, and then uniqueness and stability results can be derived [2, 23, 25].
For the elastic wave equations, the inverse problem is less understood. Rachele [18] proved that the jets of $\lambda,\mu,\rho$ at the boundary can be fully recovered. For interior determination, the boundary control method has not been successfully applied to the full Lamé equation (1). For a Lamé-type equation having the same principal part that admits a full $P/S$-decoupling, the boundary control was used to recover the wavespeeds [3, 4, 12]. Via reduction to geometric inverse problems, the uniqueness of the two wavespeeds $c_{P},c_{S}$ (or equivalently $\frac{\lambda}{\rho},\frac{\mu}{\rho}$) is known (1) assuming $(\Omega,g_{P/S})$ is simple [19]; or (2) assuming that $\partial\Omega$ is convex with respect to $g_{P/S}$ and $(\Omega,g_{P/S})$ admits a strictly convex function [26] (known as the strictly convex foliation condition, see [30, 16] for more discussions on this condition). The proof is based on recovering the lens relations related to the two wavespeeds and the applying the known rigidity results [10, 24]. With $c_{P}$ and $c_{S}$ already determined, under the above two geometric assumptions, [20, 6] study the recovery of the density $\rho$ using the subprincipal term of the $P$-wave amplitude and the inversion of the longitudinal geodesic ray transform of a second order tensor. But the results need to essentially assume a technical assumption $\lambda\neq 2\mu$. These results can be extended to uniqueness of piecewise smooth parameters [8, 7]. We also refer to [9] for a uniqueness result for the density under different assumptions. For a nonlinear elastic wave equation, the density can be recovered utilizing the nonlinearity [31].
In this work we study the recovery of the density $\rho$ using $S$-wave amplitude with known $c_{P}$ and $c_{S}$. With $c_{P/S}$ already known, we use the notation $\Lambda_{\rho}=\Lambda_{\lambda,\mu,\rho}$ for simplicity. The readers will see that the problem reduces to the inversion of the transverse geodesic ray transform of a second order tensor. The invertibility of this type of ray transform is proved recently in [32, 33] under the strictly convex foliation condition. As pointed out in [24] the foliation condition is a natural generalization of the Herglotz [13] and the Wieckert-Zoeppritz [34] conditions, which are typically assumed for the Earth. The main result of this paper is as follows.
Theorem 1.
Let $c_{P}$ and $c_{S}$ be already known. Assume $\partial\Omega$ is strictly convex with respect to $g_{S}$, and $(\Omega,g_{S})$ admits a strictly convex function. If $T$ is greater than the length of all geodesics in $(\Omega,g_{S})$, then $\Lambda_{\rho}$ determines $\rho$ in $\overline{\Omega}$ uniquely.
Remark 1.
Compared with the results [20, 6], we have removed the artificial assumption $\lambda\neq 2\mu$.
Remark 2.
Under the strictly convex foliation conditions for both $g_{P}$ and $g_{S}$, we indeed have the the uniqueness of all parameters $\lambda,\mu,\rho$, with preliminary uniqueness results for $\frac{\lambda}{\rho}$ and $\frac{\mu}{\rho}$ in [26]. To be more precise, we have the following corollary.
Corollary 1.
Assume $\partial\Omega$ is strictly convex with respect to $g_{P}$ and $g_{S}$, and $(\Omega,g_{P/S})$ admits a strictly convex function. If $T$ is greater than the length of all geodesics in $(\Omega,g_{P/S})$, then $\Lambda_{\lambda,\mu,\rho}$ determines $\lambda,\mu,\rho$ in $\overline{\Omega}$ uniquely.
Remark 3.
Under the simplicity condition, the injectivity of the related transverse ray transform is also known when the curvature is close to zero [21], and one can directly have uniqueness of $\rho$ under the same assumption for $(\Omega,g_{S})$.
Remark 4.
As in [26, 6], one can also prove a local uniqueness result, namely, for recovering the parameter near a convex part of the boundary from measurements taken only on that part.
2. Geometric optics solutions
The proof relies on the construction of geometric optics solutions expressed using Fourier integral operators [14, 11]. Assume $V$ is an open subset of $\partial\Omega$. We will construct such solutions to the elastic wave equation (1) of the form
(4) $u(t,x)=(2\pi)^{-1}\int_{\mathbb{R}}e^{\mathrm{i}\varrho(t-\varphi_{P/S}(x))}\mathbf{a}(t,x,\varrho)\mathrm{d}\varrho,$
where $\mathbf{a}\in S^{m}$ is a vector. The wavefront set of the above distribution is contained in
(5) $\{(t,x,\varrho,\varrho\nabla_{x}\varphi_{P/S});t=\varphi_{P/S}(x),\varrho\neq 0\}.$
The phase functions $\varphi_{P/S}(x)$ solve the eikonel equations
$c^{2}_{P/S}|\nabla_{x}\varphi_{P/S}|^{2}=1.$
The solution to the eikonal equation exists only locally in general. Then the wave paths are given by
$(\tau,\xi)=\varrho(1,\nabla_{x}\varphi_{P/S}).$
Fix $(t_{0},x_{0},\tau_{0},\xi_{0})$, $x_{0}\in\partial\Omega$, in the set (5). Take $V\subset\partial\Omega$ a neighborhood of $x_{0}$. We can microlocalize $u$ such that it is defined near the ($P$- or $S$-wave) null-geodesic $\vartheta_{0}$ issued from $(t_{0},x_{0})\in(0,T)\times V$ in the direction $(\tau_{0},\xi_{0})$ but in some neighborhood of $(t_{0},x_{0})$. To extend it globally, we assume that it is valid until $\vartheta_{0}$ hits another hypersurface $(0,T)\times V_{1}$, and a bit beyond it. We then consider a new similar problem with $V$ replaced by $V_{1}$. By a compactness argument, we can cover the whole null-geodesic.
Assume that $\mathbf{a}$ admits an asymptotic expansion as $\varrho\rightarrow\infty$
$\mathbf{a}(t,x,\varrho)\sim\sum_{j=0}^{\infty}\varrho^{-j}\mathbf{a}_{j}(t,x).$
We will carry out the construction of $\mathbf{a}$ in the so-called ray coordinates (associated with $g_{\bullet}=c^{-2}_{\bullet}\mathrm{d}s^{2}$) for $\mathbb{R}^{3}$, i.e., curvilinear coordinates $(x^{1},x^{2},x^{3})$ such that $x^{3}(x)=\tau=\varphi_{\bullet}(x)$ and the coordinate surface $x^{3}=x^{3}_{0}$ are orthogonal to the coordinate lines $x^{1}=x_{0}^{1}$, $x^{2}=x^{2}_{0}$ that are the geodesics of metric $g_{\bullet}$. Here $\bullet=P$ or $S$, and we occasionally omit it. Under these ray coordinates the Euclidean metric has the form
$\mathrm{d}s^{2}=g_{\alpha\beta}\mathrm{d}x^{\alpha}\mathrm{d}x^{\beta}+c^{2}_{\bullet}\mathrm{d}\tau^{2}.$
From now on the Greek indices range over the values $1,2$. The Christoffel symbols of the Euclidean metric are then
$\begin{split}&\Gamma^{3}_{\alpha\beta}=-\frac{1}{2}c^{-2}\frac{\partial g_{\alpha\beta}}{\partial\tau},\quad\Gamma^{\beta}_{\alpha 3}=\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau},\quad\Gamma^{3}_{\alpha 3}=c^{-1}\frac{\partial c}{\partial x^{\alpha}},\\ &\Gamma^{\alpha}_{33}=-cg^{\alpha\beta}\frac{\partial c}{\partial x^{\beta}},\quad\Gamma^{3}_{33}=c^{-1}\frac{\partial c}{\partial\tau}.\end{split}$
Note that the equation (1) can be written as
$\rho\partial^{2}_{t}u-\nabla\cdot\sigma(u)=0,$
where
$\varepsilon(u)=\frac{1}{2}(\nabla u+(\nabla u)^{T})$
is the strain tensor and
$\sigma(u)=\lambda\mathrm{trace}(\varepsilon(u))e+2\mu\varepsilon(u)$
is the stress tensor. Here $e$ is the Euclidean metric tensor.
For ease of notation we write $\varphi=\varphi_{P}$ or $\varphi_{P}$. Denote
$u_{\varrho}=\mathbf{a}e^{\mathrm{i}(t-\varrho\varphi)},$
and write $\mathbf{a}=(a_{1},a_{2},a_{3})$ under the ray coordinates. The strain tensor takes the form
$\varepsilon_{k\ell}(u_{\varrho})=\frac{1}{2}\mathrm{i}\varrho(a_{k}\varphi_{;\ell}+a_{\ell}\varphi_{;k})e^{\mathrm{i}\varrho\varphi}+\frac{1}{2}(a_{k;\ell}+a_{\ell;k})e^{\mathrm{i}\varrho\varphi}.$
The stress tensor is then
$\begin{split}\sigma_{ij}(u_{\varrho})=&\lambda e^{k\ell}\varepsilon_{k\ell}e_{ij}+2\mu\varepsilon_{ij}\\ =&\lambda g^{k\ell}\varepsilon_{k\ell}g_{ij}+2\mu\varepsilon_{ij}\\ =&\mathrm{i}\varrho(\lambda a_{k}\varphi_{;\ell}g^{k\ell}g_{ij}+\mu a_{i}\varphi_{;j}+\mu a_{j}\varphi_{;i})e^{\mathrm{i}\varrho\varphi}+(\lambda a_{k;\ell}g^{k\ell}g_{ij}+\mu a_{i;j}+\mu a_{j;i})e^{\mathrm{i}\varrho\varphi}.\end{split}$
We calculate
$\sigma_{ij;m}(u_{\varrho})=$ $\partial_{m}\sigma_{ij}-\Gamma^{n}_{im}\sigma_{nj}-\Gamma^{n}_{jm}\sigma_{ni}$
$=$ $-\varrho^{2}(\lambda a_{k}\varphi_{;\ell}g^{k\ell}g_{ij}+\mu a_{i}\varphi_{;j}+\mu a_{j}\varphi_{;i})\varphi_{;m}e^{\mathrm{i}\varrho\varphi}$
$+\mathrm{i}\varrho\partial_{m}(\lambda a_{k}\varphi_{;\ell}g^{k\ell}g_{ij}+\mu a_{i}\varphi_{;j}+\mu a_{j}\varphi_{;i})e^{\mathrm{i}\varrho\varphi}$
$+\mathrm{i}\varrho(\lambda a_{k;\ell}g^{k\ell}g_{ij}+\mu a_{i;j}+\mu a_{j;i})\varphi_{;m}e^{\mathrm{i}\varrho\varphi}$
$-\mathrm{i}\varrho\Gamma^{n}_{im}(\lambda a_{k}\varphi_{;\ell}g^{k\ell}g_{nj}+\mu a_{n}\varphi_{;j}+\mu a_{j}\varphi_{;n})e^{\mathrm{i}\varrho\varphi}$
$-\mathrm{i}\varrho\Gamma^{n}_{jm}(\lambda a_{k}\varphi_{;\ell}g^{k\ell}g_{ni}+\mu a_{n}\varphi_{;i}+\mu a_{i}\varphi_{;n})e^{\mathrm{i}\varrho\varphi}$
$+\partial_{m}(\lambda a_{k;\ell}g^{k\ell}g_{ij}+\mu a_{i;j}+\mu a_{j;i})e^{\mathrm{i}\varrho\varphi}$
$-\Gamma_{im}^{n}(\lambda a_{k;\ell}g^{k\ell}g_{nj}+\mu a_{n;j}+\mu a_{j;n})e^{\mathrm{i}\varrho\varphi}$
$-\Gamma_{jm}^{n}(\lambda a_{k;\ell}g^{k\ell}g_{ni}+\mu a_{n;i}+\mu a_{i;n})e^{\mathrm{i}\varrho\varphi}.$
Therefore
$\sigma_{ij;}^{~{}\,~{}j}(u_{\varrho})=$ $\sigma_{ij;m}g^{jm}$
$=$ $-\varrho^{2}(\lambda a_{k}\varphi_{;\ell}g^{k\ell}\varphi_{;i}+\mu a_{j}\varphi_{;m}g^{jm}\varphi_{;i}+\mu\varphi_{;j}\varphi_{;m}g^{jm}a_{i})e^{\mathrm{i}\varrho\varphi}$
$+\mathrm{i}\varrho\left(\partial_{i}(\lambda a_{k}\varphi_{;\ell}g^{k\ell})+\lambda a_{k}\varphi_{;\ell}g^{k\ell}\partial_{m}g_{ij}g^{jm}+\partial_{m}(\mu a_{i}\varphi_{;j}+\mu a_{j}\varphi_{;i})g^{jm}\right)e^{\mathrm{i}\varrho\varphi}$
$+\mathrm{i}\varrho\left(\lambda a_{k;\ell}g^{k\ell}\varphi_{;i}+\mu(a_{i;j}+a_{j;i})\varphi_{;m}g^{jm}\right)e^{\mathrm{i}\varrho\varphi}$
$-\mathrm{i}\varrho\Gamma^{n}_{im}(\lambda a_{k}\varphi_{;\ell}g^{k\ell}g_{nj}+\mu a_{n}\varphi_{;j}+\mu a_{j}\varphi_{;n})g^{jm}e^{\mathrm{i}\varrho\varphi}$
$-\mathrm{i}\varrho\Gamma^{n}_{jm}(\lambda a_{k}\varphi_{;\ell}g^{k\ell}g_{ni}+\mu a_{n}\varphi_{;i}+\mu a_{i}\varphi_{;n})g^{jm}e^{\mathrm{i}\varrho\varphi}$
$+\partial_{m}(\lambda a_{k;\ell}g^{k\ell}g_{ij}+\mu a_{i;j}+\mu a_{j;i})g^{jm}e^{\mathrm{i}\varrho\varphi}$
$-\Gamma_{im}^{n}(\lambda a_{k;\ell}g^{k\ell}g_{nj}+\mu a_{n;j}+\mu a_{j;n})g^{jm}e^{\mathrm{i}\varrho\varphi}$
$-\Gamma_{jm}^{n}(\lambda a_{k;\ell}g^{k\ell}g_{ni}+\mu a_{n;i}+\mu a_{i;n})g^{jm}e^{\mathrm{i}\varrho\varphi}.$
[…]
$\rho\partial^{2}_{t}u_{\varrho}=-\varrho^{2}\rho\mathbf{a}e^{\mathrm{i}\varrho(t-\varphi)}+2\mathrm{i}\varrho\rho(\partial_{t}\mathbf{a})e^{\mathrm{i}\varrho(t-\varphi)}+\rho\partial^{2}_{t}\mathbf{a}e^{\mathrm{i}\varrho(t-\varphi)}.$
In order to construct microlocal solutions to (1), we need to take $u_{\varrho}$ such that
(6) $\rho\partial^{2}_{t}(u_{\varrho})_{i}-\sigma_{ij;}^{~{}\,~{}j}(u_{\varrho})\sim 0.$
One can use the above equation to derive a sequence of equations for $\varphi$ and $\mathbf{a}_{j}$ by equating the coefficients of same orders in $\varrho$. We refer to [19, 20] for a construction of microlocal solutions under Cartesian corrdinates.
Equating the coefficients of order $\varrho^{2}$ in the equation (6), we get
(7) $\rho\mathbf{a}_{0}=(\lambda+\mu)\langle\mathbf{a}_{0},\nabla\varphi\rangle\nabla\varphi+\mu|\nabla\varphi|^{2}\mathbf{a}_{0}.$
For $P$ waves, the wave vector $\nabla\varphi$ and the principal amplitude/polarization $\mathbf{a}_{0}$ are parallel; for $S$-waves, they are orthogonal. We will use the subprincipal polarization $\mathbf{a}_{1}$ to recover the density. The explicit construction of $\varphi$, $\mathbf{a}_{0}$ and $\mathbf{a}_{1}$ will be carried out in the following two sections.
3. P-waves and related results
In this section, we first construct solutions of the form (4) that represent $P$-waves. We will regain the same results obtained in [20, 6], but the calculation is done in a different manner (under curvilinear coordinates versus Cartesian coordinates). For $P$-waves, we take $\varphi=\varphi_{P}$ such that
$|\nabla\varphi|^{2}=c_{P}^{-2},$
with $\nabla\varphi(x_{0})=\xi_{0}$. In order for (7) to hold, we can take
$\mathbf{a}_{0}=A_{P}\frac{\nabla\varphi}{|\nabla\varphi|},$
where $A_{P}$ is a scalar function. Under the ray coordinates associated with $g_{P}$,
$\varphi_{;3}=1,\varphi_{;}^{~{}~{}3}=c_{P}^{-2},\varphi_{;\alpha}=\varphi_{;}^{~{}~{}\alpha}=0.$
Therefore
$a_{03}=c_{P}A_{P}\quad a_{0}^{~{}3}=c_{P}^{-1}A_{P},\quad a_{0\alpha}=a_{0}^{~{}\alpha}=0,\,\alpha=1,2.$
To determine $A_{P}$, we calculate
$\partial_{3}(\lambda a_{0k}\varphi_{;\ell}g^{k\ell})+\lambda a_{0k}\varphi_{;\ell}g^{k\ell}\partial_{m}g_{3j}g^{jm}+\partial_{m}(\mu a_{03}\varphi_{;j}+\mu a_{0j}\varphi_{;3})g^{jm}$
$=\partial_{3}(\lambda c_{P}A_{P}c_{P}^{-2})+\lambda c_{P}A_{P}c_{P}^{-2}\partial_{3}g_{33}g^{33}+\partial_{3}(\mu c_{P}A_{P}+\mu c_{P}A_{P})g^{33}$
$=\partial_{3}(\lambda c_{P}^{-1}A_{P})+\lambda c_{P}^{-3}A_{P}\partial_{3}(c_{P}^{2})+\partial_{3}(2\mu c_{P}A_{P})c_{P}^{-2}$
$=\partial_{3}((\lambda+2\mu)c_{P}A_{P})c_{P}^{-2},$
$\lambda a_{0k;\ell}g^{k\ell}\varphi_{;3}+\mu(a_{03;j}+a_{0j;3})\varphi_{;m}g^{jm}$
$=$ $\lambda(\partial_{3}(c_{P}A_{P})c_{P}^{-2}-c_{P}A_{P}\Gamma_{33}^{3}c_{P}^{-2}-c_{P}A_{P}\Gamma^{3}_{\alpha\beta}g^{\alpha\beta})+\mu(\partial_{3}(c_{P}A_{P})-c_{P}A_{P}\Gamma_{33}^{3}+\partial_{3}(c_{P}A_{P})-c_{P}A_{P}\Gamma_{33}^{3})c^{-2}_{P}$
$=$ $(\lambda+2\mu)\partial_{3}(c_{P}A_{P})c_{P}^{-2}-(\lambda+2\mu)A_{P}(\partial_{3}c_{P})c_{P}^{-2}+\frac{1}{2}\lambda c_{P}^{-1}A_{P}\frac{\partial g_{\alpha\beta}}{\partial\tau}g^{\alpha\beta},$
$\Gamma^{n}_{3m}(\lambda a_{0k}\varphi_{;\ell}g^{k\ell}g_{nj}+\mu a_{0n}\varphi_{;j}+\mu a_{0j}\varphi_{;n})g^{jm}$
$=$ $\Gamma_{3n}^{n}\lambda c_{P}A_{P}g^{33}+\Gamma_{33}^{3}(\mu c_{P}A_{P}+\mu c_{P}A_{P})g^{33}$
$=$ $\frac{1}{2}g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}\lambda c_{P}^{-1}A_{P}+c_{P}^{-2}\partial_{3}c_{P}(\lambda+2\mu)A_{P},$
$\Gamma^{n}_{jm}(\lambda a_{0k}\varphi_{;\ell}g^{k\ell}g_{n3}+\mu a_{0n}\varphi_{;3}+\mu a_{03}\varphi_{;n})g^{jm}$
$=$ $\Gamma_{jm}^{3}(\lambda a_{03}g^{33}g_{33}+\mu a_{03}+\mu a_{03})g^{jm}$
$=$ $(c_{P}^{-3}\partial_{3}c_{P}-\frac{1}{2}c_{P}^{-2}\frac{\partial g_{\alpha\beta}}{\partial\tau}g^{\alpha\beta})(\lambda+2\mu)c_{P}A_{P}.$
Equating the coefficients of order $\mathcal{O}(\varrho)$ in the equation (6) with $i=3$, one obtains
$\begin{split}&-(\lambda a_{13}c_{P}^{-2}+2\mu a_{13}c_{P}^{-2}-\rho a_{13})\\ &+\mathrm{i}c_{P}^{-1}\left(2(\lambda+2\mu)(\partial_{3}-\partial_{t})A_{P}-(\lambda+2\mu)c_{P}^{-1}(\partial_{3}c_{P})A_{P}+\partial_{3}(\lambda+2\mu)A_{P}+\frac{1}{2}(\lambda+2\mu)\frac{\partial g_{\alpha\beta}}{\partial\tau}g^{\alpha\beta}\right)=0.\end{split}$
We can take $A_{P}$ to be independent of $t$. Using the equality
$g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}=\frac{2}{J}\frac{\partial J}{\partial\tau},$
where $J^{2}=\det(g_{\alpha\beta})$, we end up with
$J(\lambda+2\mu)\partial_{\tau}A_{P}-\frac{1}{2}J(\lambda+2\mu)c_{P}^{-1}(\partial_{\tau}c_{P})A_{P}+\frac{1}{2}J\partial_{\tau}(\lambda+2\mu)A_{P}+\frac{1}{2}(\partial_{\tau}J)(\lambda+2\mu)A_{P}=0.$
The above equation can be considered as a linear ODE along a single geodesic $\gamma$. Note that it can be written as
$\frac{\partial}{\partial\tau}(A_{P}J^{1/2}c_{P}^{-1/2}(\lambda+2\mu)^{1/2})=0.$
Solving the above equation, one gets
$A_{P}=cJ^{-1/2}c_{P}^{1/2}(\lambda+2\mu)^{-1/2}=cJ^{-1/2}c_{P}^{-1/2}\rho^{-1/2}$
with an arbitrary constant $c$. Since we need to use nontrivial solution, for simplicity, we just take $c=1$.
Next, we collect the coefficients of order $\mathcal{O}(\varrho)$ in the equation (6) with $i=\alpha$. For this, we calculate
$\partial_{\alpha}(\lambda a_{0k}\varphi_{;\ell}g^{k\ell})+\lambda a_{0k}\varphi_{;\ell}g^{k\ell}\partial_{m}g_{\alpha j}g^{jm}+\partial_{m}(\mu a_{0\alpha}\varphi_{;j}+\mu a_{0j}\varphi_{;\alpha})g^{jm}$
$=$ $\partial_{\alpha}(\lambda c_{P}A_{P}c_{P}^{-2})+\lambda c_{P}A_{P}c_{P}^{-2}(\partial_{\gamma}g_{\alpha\beta})g^{\beta\gamma},$
$\lambda a_{0k;\ell}g^{k\ell}\varphi_{;\alpha}+\mu(a_{0\alpha;j}+a_{0j;\alpha})\varphi_{;m}g^{jm}$
$=$ $\mu(-\Gamma_{\alpha 3}^{3}a_{03}+\partial_{\alpha}(a_{03})-\Gamma_{\alpha 3}^{3}a_{03})g^{33}$
$=$ $\mu(\partial_{\alpha}(c_{P}A_{P})-2c_{P}^{-1}(\partial_{\alpha}c_{P})c_{P}A_{P})c_{P}^{-2},$
$\Gamma^{n}_{\alpha m}(\lambda a_{0k}\varphi_{;\ell}g^{k\ell}g_{nj}+\mu a_{0n}\varphi_{;j}+\mu a_{0j}\varphi_{;n})g^{jm}$
$=$ $\Gamma_{\alpha n}^{n}(\lambda c_{P}A_{P}c_{P}^{-2})+\Gamma_{\alpha 3}^{3}g^{33}(2\mu c_{P}A_{P})$
$=$ $\Gamma_{\alpha\beta}^{\beta}\lambda c_{P}^{-1}A_{P}+\Gamma_{\alpha 3}^{3}(\lambda+2\mu)c_{P}^{-1}A_{P},$
$\Gamma^{n}_{jm}(\lambda a_{0k}\varphi_{;\ell}g^{k\ell}g_{n\alpha}+\mu a_{0n}\varphi_{;\alpha}+\mu a_{0\alpha}\varphi_{;n})g^{jm}$
$=$ $\Gamma_{jm}^{\beta}(\lambda a_{03}g^{33}g_{\alpha\beta})g^{jm}$
$=$ $\Gamma^{\beta}_{33}\lambda c_{P}^{-1}A_{P}g_{\alpha\beta}c_{P}^{-2}+\Gamma^{\beta}_{\gamma\delta}\lambda c_{P}^{-1}A_{P}g_{\alpha\beta}g^{\gamma\delta}.$
Using the identity
(8) $(\partial_{\gamma}g_{\alpha\beta})g^{\beta\gamma}-\Gamma_{\alpha\beta}^{\beta}-\Gamma^{\beta}_{\gamma\delta}g_{\alpha\beta}g^{\gamma\delta}=0,$
we obtain
$\begin{split}&-\mu c_{P}^{-2}a_{1\alpha}+\rho a_{1\alpha}\\ &+\mathrm{i}(\partial_{\alpha}(\lambda A_{P})+\mu(\partial_{\alpha}(c_{P}A_{P})-2(\partial_{\alpha}c_{P})A_{P})c_{P}^{-2}-\Gamma_{\alpha 3}^{3}(\lambda+2\mu)c_{P}^{-1}A_{P}-\Gamma^{\beta}_{33}\lambda c_{P}^{-3}A_{P}g_{\alpha\beta})=0,\end{split}$
which simplifies to
$-\mu c_{P}^{-2}a_{1\alpha}+\rho a_{1\alpha}+\mathrm{i}((\lambda+\mu)c_{P}^{-1}\partial_{\alpha}A_{P}+(\partial_{\alpha}\lambda)c_{P}^{-1}A_{P}-c_{P}^{-2}\partial_{\alpha}c_{P}(3\mu+\lambda)A_{P})=0.$
So, for $\alpha=1,2$,
(9) $\begin{split}a_{1\alpha}=&\frac{\mathrm{i}}{c_{S}^{2}c_{P}^{-2}-1}\left(\frac{\lambda+\mu}{\rho}c_{P}^{-1}\partial_{\alpha}A_{P}+\frac{(\partial_{\alpha}\lambda)}{\rho}c_{P}^{-1}A_{P}-c_{P}^{-2}\partial_{\alpha}c_{P}\frac{3\mu+\lambda}{\rho}A_{P}\right)\\ =&\frac{\mathrm{i}}{c_{S}^{2}c_{P}^{-2}-1}\Big{(}(c_{P}^{2}-c_{S}^{2})c_{P}^{-1}J^{-1/2}c_{P}^{-1/2}(-\frac{1}{2}\rho^{-3/2}\partial_{\alpha}\rho)+(c_{P}^{2}-c_{S}^{2})c_{P}^{-1}\partial_{\alpha}(J^{-1/2}c_{P}^{-1/2})\rho^{-1/2}\\ &+\rho^{-3/2}(c_{P}^{2}-2c_{S}^{2})(\partial_{\alpha}\rho)c_{P}^{-1}J^{-1/2}c_{P}^{-1/2}+\partial_{\alpha}(c_{P}^{2}-2c_{S}^{2})c_{P}^{-1}J^{-1/2}c_{P}^{-1/2}\rho^{-1/2}\\ &-c_{P}^{-2}(\partial_{\alpha}c_{P})(c_{P}^{2}+c_{S}^{2})J^{-1/2}c_{P}^{-1/2}\rho^{-1/2}\Big{)}\\ =&\frac{\mathrm{i}}{c_{S}^{2}c_{P}^{-2}-1}\Big{(}\frac{1}{2}\rho^{-3/2}(c_{P}^{2}-3c_{S}^{2})(\partial_{\alpha}\rho)c_{P}^{-3/2}J^{-1/2}+(c_{P}^{2}-c_{S}^{2})c_{P}^{-3/2}\partial_{\alpha}(J^{-1/2})\rho^{-1/2}\\ &\quad\quad\quad\quad+c_{P}^{-5/2}\left(\frac{1}{2}c_{P}^{2}\partial_{\alpha}c_{P}-\frac{1}{2}c_{S}^{2}\partial_{\alpha}c_{P}-4c_{S}c_{P}\partial_{\alpha}c_{S}\right)J^{-1/2}\rho^{-1/2}\Big{)}\\ =&\frac{\mathrm{i}}{(c_{S}^{2}c_{P}^{-2}-1)\sqrt{\rho}}\Big{(}(c_{P}^{2}-3c_{S}^{2})(\partial_{\alpha}\log\sqrt{\rho})c_{P}^{-3/2}J^{-1/2}+(c_{P}^{2}-c_{S}^{2})c_{P}^{-3/2}\partial_{\alpha}(J^{-1/2})\\ &\quad\quad\quad\quad+c_{P}^{-5/2}\left(\frac{1}{2}c_{P}^{2}\partial_{\alpha}c_{P}-\frac{1}{2}c_{S}^{2}\partial_{\alpha}c_{P}-4c_{S}c_{P}\partial_{\alpha}c_{S}\right)J^{-1/2}\Big{)}\\ =&\frac{\mathrm{i}}{(c_{S}^{2}-c_{P}^{2})\sqrt{\rho}}\Big{(}(c_{P}^{2}-3c_{S}^{2})(\partial_{\alpha}\log\sqrt{\rho})c_{P}^{1/2}J^{-1/2}+(c_{P}^{2}-c_{S}^{2})c_{P}^{1/2}\partial_{\alpha}(J^{-1/2})\\ &\quad\quad\quad\quad+c_{P}^{-1/2}\left(\frac{1}{2}c_{P}^{2}\partial_{\alpha}c_{P}-\frac{1}{2}c_{S}^{2}\partial_{\alpha}c_{P}-4c_{S}c_{P}\partial_{\alpha}c_{S}\right)J^{-1/2}\Big{)}.\end{split}$
We need to use the $\mathcal{O}(1)$ terms in the equation (6) with $i=3$, that is,
$\begin{split}&2\mathrm{i}\rho\partial_{t}a_{13}\\ =&\mathrm{i}\left(\partial_{3}(\lambda a_{1k}\varphi_{;\ell}g^{k\ell})+\lambda a_{1k}\varphi_{;\ell}g^{k\ell}\partial_{m}g_{3j}g^{jm}+\partial_{m}(\mu a_{13}\varphi_{;j}+\mu a_{1j}\varphi_{;3})g^{jm}\right)\\ &\mathrm{i}\left(\lambda a_{1k;\ell}g^{k\ell}\varphi_{;3}+\mu(a_{13;j}+a_{1j;3})\varphi_{;m}g^{jm}\right)\\ &-\mathrm{i}\Gamma^{n}_{3m}(\lambda a_{1k}\varphi_{;\ell}g^{k\ell}g_{nj}+\mu a_{1n}\varphi_{;j}+\mu a_{1j}\varphi_{;n})g^{jm}\\ &-\mathrm{i}\Gamma^{n}_{jm}(\lambda a_{1k}\varphi_{;\ell}g^{k\ell}g_{n3}+\mu a_{1n}\varphi_{;3}+\mu a_{13}\varphi_{;n})g^{jm}\\ &+\partial_{m}(\lambda a_{0k;\ell}g^{k\ell}g_{3j}+\mu a_{03;j}+\mu a_{0j;3})g^{jm}\\ &-\Gamma_{3m}^{n}(\lambda a_{0k;\ell}g^{k\ell}g_{nj}+\mu a_{0n;j}+\mu a_{0j;n})g^{jm}\\ &-\Gamma_{jm}^{n}(\lambda a_{0k;\ell}g^{k\ell}g_{n3}+\mu a_{0n;3}+\mu a_{3;n})g^{jm}.\end{split}$
Then we calculate
$\partial_{\beta}a_{1\alpha}=$ $\frac{\mathrm{i}}{(c_{S}^{2}-c_{P}^{2})\sqrt{\rho}}\Bigg{(}(c_{P}^{2}-3c_{S}^{2})(\partial_{\beta}\partial_{\alpha}\log\sqrt{\rho})c_{P}^{1/2}J^{-1/2}+\partial_{\beta}((c_{P}^{2}-3c_{S}^{2})c_{P}^{1/2}J^{-1/2})\partial_{\alpha}\log\sqrt{\rho}$
$\quad\quad\quad\quad\quad-\frac{\partial_{\beta}c_{S}^{2}-\partial_{\beta}c_{P}^{2}}{(c_{S}^{2}-c_{P}^{2})}(c_{P}^{2}-3c_{S}^{2})(\partial_{\alpha}\log\sqrt{\rho})c_{P}^{1/2}J^{-1/2}$
$\quad\quad\quad\quad\quad+(c_{P}^{2}-3c_{S}^{2})(\partial_{\alpha}\log\sqrt{\rho})c_{P}^{1/2}J^{-1/2}(-\partial_{\beta}\log\sqrt{\rho})$
$\quad\quad\quad\quad\quad+(c_{P}^{2}-c_{S}^{2})c_{P}^{1/2}\partial_{\alpha}(J^{-1/2})(-\partial_{\beta}\log\sqrt{\rho})$
$\quad\quad\quad\quad\quad+c_{P}^{-1/2}\left(\frac{1}{2}c_{P}^{2}\partial_{\alpha}c_{P}-\frac{1}{2}c_{S}^{2}\partial_{\alpha}c_{P}-4c_{S}c_{P}\partial_{\alpha}c_{S}\right)J^{-1/2}(-\partial_{\beta}\log\sqrt{\rho})+C\Bigg{)}.$
With the above expression, we get
$\partial_{3}(\lambda a_{1k}\varphi_{;\ell}g^{k\ell})+\lambda a_{1k}\varphi_{;\ell}g^{k\ell}\partial_{m}g_{3j}g^{jm}+\partial_{m}(\mu a_{13}\varphi_{;j}+\mu a_{1j}\varphi_{;3})g^{jm}$
$=$ $\partial_{3}(\lambda a_{13}g^{33})+\lambda a_{13}g^{33}\partial_{3}(g_{33})g^{33}+\partial_{3}(\mu a_{13})g^{33}+\partial_{m}(\mu a_{1j})g^{jm}$
$=$ $\partial_{3}(\lambda a_{13}c_{P}^{-2})+\lambda a_{13}c_{P}^{-4}\partial_{3}(c_{P}^{2})+2\partial_{3}(\mu a_{13})c_{P}^{-2}+\partial_{\alpha}(\mu a_{1\beta})g^{\alpha\beta}$
$=$ $\partial_{3}((\lambda+2\mu)a_{13})c_{P}^{-2}+\partial_{\alpha}(\mu a_{1\beta})g^{\alpha\beta},$
$\lambda a_{1k;\ell}g^{k\ell}\varphi_{;3}+\mu(a_{13;j}+a_{1j;3})\varphi_{;m}g^{jm}$
$=$ $\lambda(a_{13;3}g^{33}+a_{1\alpha;\beta}g^{\alpha\beta})+2\mu a_{13;3}g^{33}$
$=$ $\lambda(\partial_{3}a_{13}-\Gamma^{j}_{33}a_{1j})c_{P}^{-2}+\lambda(\partial_{\beta}a_{1\alpha}-\Gamma^{j}_{\alpha\beta}a_{1j})g^{\alpha\beta}+2\mu(\partial_{3}a_{13}-\Gamma_{33}^{j}a_{1j})c_{P}^{-2}$
$=$ $(\lambda+2\mu)c_{P}^{-2}\partial_{3}a_{13}-(\lambda+2\mu)c_{P}^{-1}(\partial_{3}c_{P})c_{P}^{-2}a_{13}-(\lambda+2\mu)(\Gamma_{33}^{\alpha}a_{1\alpha})c_{P}^{-2}+\lambda g^{\alpha\beta}\partial_{\beta}a_{1\alpha}$
$+\frac{1}{2}\lambda c_{P}^{-2}a_{13}\frac{\partial g_{\alpha\beta}}{\partial\tau}g^{\alpha\beta}-\lambda g^{\alpha\beta}\Gamma_{\alpha\beta}^{\gamma}a_{1\gamma},$
$\Gamma^{n}_{3m}(\lambda a_{1k}\varphi_{;\ell}g^{k\ell}g_{nj}+\mu a_{1n}\varphi_{;j}+\mu a_{1j}\varphi_{;n})g^{jm}$
$=$ $\Gamma_{3n}^{n}\lambda a_{13}g^{33}+\Gamma_{33}^{n}\mu a_{1n}g^{33}+\Gamma_{3m}^{3}\mu a_{1j}g^{jm}$
$=$ $\frac{1}{2}g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}\lambda c_{P}^{-2}a_{13}+c_{P}^{-3}\partial_{3}c_{P}(\lambda+2\mu)a_{13}+\Gamma_{33}^{\alpha}\mu a_{1\alpha}c_{P}^{-2}+\Gamma_{3\beta}^{3}\mu a_{1\alpha}g^{\alpha\beta},$
$\Gamma^{n}_{jm}(\lambda a_{1k}\varphi_{;\ell}g^{k\ell}g_{n3}+\mu a_{1n}\varphi_{;3}+\mu a_{13}\varphi_{;n})g^{jm}$
$=$ $\Gamma_{jm}^{3}(\lambda a_{13}g^{33}g_{33}+\mu a_{13}+\mu a_{13})g^{jm}+\Gamma_{jm}^{\alpha}(\mu a_{1\alpha})g^{jm}$
$=$ $(c_{P}^{-3}\partial_{3}c_{P}-\frac{1}{2}c_{P}^{-2}\frac{\partial g_{\alpha\beta}}{\partial\tau}g^{\alpha\beta})(\lambda+2\mu)a_{13}+\Gamma^{\alpha}_{\beta\gamma}\mu a_{1\alpha}g^{\beta\gamma}+\Gamma^{\alpha}_{33}\mu a_{1\alpha}c_{P}^{-2}.$
We also calculate
$\partial_{m}(\lambda a_{0k;\ell}g^{k\ell}g_{3j}+\mu a_{03;j}+\mu a_{0j;3})g^{jm}$
$=$ $\partial_{3}(\lambda(\partial_{3}a_{03}-\Gamma^{j}_{33}a_{j})g^{33}g_{33}+\lambda(-\Gamma^{j}_{\alpha\beta}a_{j})g^{\alpha\beta}g_{33})g^{33}$
$\quad\quad\quad+\partial_{m}(\mu\partial_{j}a_{03}-\mu\Gamma^{i}_{3j}a_{0i})g^{jm}+\partial_{m}(\mu\partial_{3}a_{0j}-\mu\Gamma_{3j}^{i}a_{0i})g^{jm}$
$=$ $\partial_{3}(\lambda\partial_{3}(c_{P}A_{P})-\lambda\Gamma^{3}_{33}c_{P}A_{P}-\lambda\Gamma^{3}_{\alpha\beta}c_{P}A_{P}g^{\alpha\beta}c_{P}^{2})c_{P}^{-2}$
$+g^{jm}\partial_{m}(\mu\partial_{j}(c_{P}A_{P}))+g^{33}\partial_{3}(\mu\partial_{3}(c_{P}A_{P}))-2g^{jm}\partial_{m}(\mu\Gamma^{3}_{3j}c_{P}A_{P})$
$=$ $\partial_{3}\left(\lambda\partial_{3}(c_{P}A_{P})-\lambda(\partial_{3}c_{P})A_{P}+\frac{1}{2}\lambda c_{P}g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}A_{P}\right)c_{P}^{-2}+2c_{P}^{-2}\partial_{3}(\mu\partial_{3}(c_{P}A_{P}))+g^{\alpha\beta}\partial_{\beta}(\mu\partial_{\alpha}(c_{P}A_{P}))$
$-2c_{P}^{-2}\partial_{3}(\mu(\partial_{3}c_{P})A_{P})-2g^{\alpha\beta}\partial_{\alpha}(\mu(\partial_{\beta}c_{P})A_{P})$
$=$ $\partial_{3}((\lambda+2\mu)c_{P}\partial_{3}A_{P})c_{P}^{-2}+\frac{1}{2}c_{P}^{-1}\lambda g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}\partial_{3}A_{P}+g^{\alpha\beta}\partial_{\beta}(\mu\partial_{\alpha}(c_{P}A_{P}))+\frac{1}{2}\partial_{3}\lambda c_{P}^{-1}g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}A_{P}$
$-2g^{\alpha\beta}\mu(\partial_{\beta}c_{P})\partial_{\alpha}A_{P}-2g^{\alpha\beta}\partial_{\alpha}\mu\partial_{\beta}c_{P}A_{P}+\sqrt{\rho}\times\text{terms that do not depend on $\rho$},$
$\Gamma_{3m}^{n}(\lambda a_{0k;\ell}g^{k\ell}g_{nj}+\mu a_{0n;j}+\mu a_{0j;n})g^{jm}$
$=$ $\Gamma_{3n}^{n}(\lambda(\partial_{3}a_{03}-\Gamma^{j}_{33}a_{j})g^{33}-\lambda\Gamma^{j}_{\alpha\beta}a_{j}g^{\alpha\beta})+\Gamma_{3m}^{n}(\mu\partial_{j}a_{0n}+\mu\partial_{n}a_{0j}-2\mu\Gamma^{i}_{nj}a_{0i})g^{jm}$
$=$ $\Gamma_{3n}^{n}(\lambda\partial_{3}(c_{P}A_{P})g^{33})+\Gamma^{3}_{3m}(\mu\partial_{j}(c_{P}A_{P}))g^{jm}+\Gamma^{n}_{33}(\mu\partial_{n}(c_{P}A_{P}))c_{P}^{-2}$
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}$
$=$ $\Gamma_{33}^{3}(\lambda+2\mu)c_{P}\partial_{3}A_{P}c_{P}^{-2}+\Gamma_{3\alpha}^{\alpha}\lambda c_{P}\partial_{3}A_{P}c_{P}^{-2}+\Gamma_{3\alpha}^{3}g^{\alpha\beta}\mu c_{P}\partial_{\beta}A_{P}+\Gamma_{33}^{\alpha}\mu c_{P}\partial_{\alpha}A_{P}$
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}$
$=$ $c_{P}^{-2}(\partial_{3}c_{P})(\lambda+2\mu)(\partial_{3}A_{P})+\frac{1}{2}g^{\delta\gamma}\frac{\partial g_{\delta\gamma}}{\partial\tau}\lambda c_{P}^{-1}(\partial_{3}A_{P})+\sqrt{\rho}\times\text{terms that do notdepend on $\rho$},$
$\Gamma_{jm}^{n}(\lambda a_{0k;\ell}g^{k\ell}g_{n3}+\mu a_{0n;3}+\mu a_{03;n})g^{jm}$
$=$ $g^{jm}\Gamma^{3}_{jm}(\lambda+2\mu)\partial_{3}(c_{P}A_{P})+g^{jm}\Gamma^{\alpha}_{jm}\mu\partial_{\alpha}(c_{P}A_{P})+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}$
$=$ $c_{P}^{-2}c_{P}^{-1}(\partial_{3}c_{P})(\lambda+2\mu)c_{P}\partial_{3}A_{P}+g^{\alpha\beta}(-\frac{1}{2}c_{P}^{-2}\frac{\partial g_{\alpha\beta}}{\partial\tau})(\lambda+2\mu)c_{P}\partial_{3}A_{P}$
$+c_{P}^{-2}(-c_{P}g^{\alpha\beta}(\partial_{\beta}c_{P}))\mu c_{P}\partial_{\alpha}A_{P}+g^{\gamma\delta}\Gamma^{\alpha}_{\gamma\delta}\mu c_{P}\partial_{\alpha}A_{P}+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}$
$=$ $c_{P}^{-2}(\partial_{3}c_{P})(\lambda+2\mu)(\partial_{3}A_{P})-\frac{1}{2}g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}(\lambda+2\mu)c_{P}^{-1}(\partial_{3}A_{P})$
$-g^{\alpha\beta}(\partial_{\beta}c_{P})\mu(\partial_{\alpha}A_{P})+g^{\gamma\delta}\Gamma^{\alpha}_{\gamma\delta}\mu c_{P}(\partial_{\alpha}A_{P})+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.$
We end up with an equation for $a_{13}$, which is taken to be independent of $t$,
(10) $\begin{split}&2\mathrm{i}(\lambda+2\mu)c_{P}^{-2}\partial_{3}a_{13}-3\mathrm{i}(\lambda+2\mu)c_{P}^{-3}\partial_{3}c_{P}a_{13}+\mathrm{i}\partial_{3}(\lambda+2\mu)c_{P}^{-2}a_{13}+\frac{1}{2}\mathrm{i}c_{P}^{-2}g^{\alpha\beta}\frac{g_{\alpha\beta}}{\partial\tau}(\lambda+2\mu)a_{13}\\ &+\mathrm{i}H_{1}+H_{2}=0,\end{split}$
where
$H_{1}=(\lambda+\mu)g^{\alpha\beta}\partial_{\alpha}a_{1\beta}+g^{\alpha\beta}\partial_{\alpha}\mu a_{1\beta}-(\lambda+4\mu)c_{P}^{-2}\Gamma^{\alpha}_{33}a_{1\alpha}-(\lambda+\mu)g^{\alpha\beta}\Gamma^{\gamma}_{\alpha\beta}a_{1\gamma}-\mu c_{P}^{-2}\Gamma^{\alpha}_{33}a_{1\alpha},$
and
$\begin{split}H_{2}=&(\lambda+2\mu)c_{P}^{-1}\partial^{2}_{33}A_{P}+\partial_{3}(\lambda+2\mu)c_{P}^{-1}\partial_{3}A_{P}-c_{P}^{-2}(\partial_{3}c_{P})(\lambda+2\mu)(\partial_{3}A_{P})+g^{\alpha\beta}\partial_{\beta}(\mu\partial_{\alpha}(c_{P}A_{P}))\\ &+\frac{1}{2}\partial_{3}\lambda c_{P}^{-1}g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}A_{P}-2g^{\alpha\beta}\partial_{\alpha}\mu\partial_{\beta}c_{P}A_{P}\\ &-g^{\alpha\beta}\mu\partial_{\beta}c_{P}\partial_{\alpha}A_{P}+\frac{1}{2}g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}(\lambda+2\mu)c_{P}^{-1}(\partial_{3}A_{P})-g^{\gamma\delta}\Gamma^{\alpha}_{\gamma\delta}\mu c_{P}(\partial_{\alpha}A_{P})\\ &+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.\end{split}$
Inserting the expression of $a_{1\alpha}$ (9), we obtain
$H_{1}=\frac{\mathrm{i}\sqrt{\rho}}{(c_{S}^{2}-c_{P}^{2})}\Bigg{(}(c_{P}^{2}-3c_{S}^{2})g^{\alpha\beta}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}c_{P}^{1/2}J^{-1/2}+g^{\alpha\beta}\partial_{\beta}\left((c_{P}^{2}-3c_{S}^{2})c_{P}^{1/2}J^{-1/2})\right)\partial_{\alpha}\log\sqrt{\rho}$
$-\frac{\partial_{\beta}c_{S}^{2}-\partial_{\beta}c_{P}^{2}}{(c_{S}^{2}-c_{P}^{2})}(c_{P}^{2}-3c_{S}^{2})(\partial_{\alpha}\log\sqrt{\rho})c_{P}^{1/2}J^{-1/2}$
$-(c_{P}^{2}-3c_{S}^{2})g^{\alpha\beta}(\partial_{\alpha}\log\sqrt{\rho})(\partial_{\beta}\log\sqrt{\rho})c_{P}^{1/2}J^{-1/2}-(c_{P}^{2}-c_{S}^{2})c_{P}^{1/2}g^{\alpha\beta}\partial_{\alpha}(J^{-1/2})\partial_{\beta}\log\sqrt{\rho}$
$-\frac{1}{2}c_{P}^{-1/2}(c_{P}^{2}-c_{S}^{2})g^{\alpha\beta}(\partial_{\alpha}c_{P})(\partial_{\beta}\log\sqrt{\rho})J^{-1/2}+4c_{S}c_{P}^{1/2}g^{\alpha\beta}(\partial_{\alpha}c_{S})J^{-1/2}(\partial_{\beta}\log\sqrt{\rho})\Bigg{)}$
$+g^{\alpha\beta}(2c_{S}^{2}\rho\partial_{\alpha}\log\sqrt{\rho}+2c_{S}\partial_{\alpha}c_{S}\rho)\frac{\mathrm{i}}{(c_{S}^{2}-c_{P}^{2})\sqrt{\rho}}\left((c_{P}^{2}-3c_{S}^{2})(\partial_{\beta}\log\sqrt{\rho})c_{P}^{1/2}J^{-1/2}\right)$
$+2g^{\alpha\beta}c_{S}^{2}\rho\partial_{\alpha}\log\sqrt{\rho}\frac{\mathrm{i}}{(c_{S}^{2}-c_{P}^{2})\sqrt{\rho}}(c_{P}^{2}-c_{S}^{2})c_{P}^{1/2}\partial_{\beta}(J^{-1/2})$
$+2g^{\alpha\beta}c_{S}^{2}\rho\partial_{\beta}\log\sqrt{\rho}\frac{\mathrm{i}}{(c_{S}^{2}-c_{P}^{2})\sqrt{\rho}}c_{P}^{-1/2}\left(\frac{1}{2}c_{P}^{2}\partial_{\alpha}c_{P}-\frac{1}{2}c_{S}^{2}\partial_{\alpha}c_{P}-4c_{S}c_{P}\partial_{\alpha}c_{S}\right)J^{-1/2}$
$-(\lambda+4\mu)(-c_{P}g^{\alpha\beta}\partial_{\beta}c_{P})c_{P}^{-2}\frac{\mathrm{i}}{(c_{S}^{2}-c_{P}^{2})\sqrt{\rho}}\left((c_{P}^{2}-3c_{S}^{2})(\partial_{\alpha}\log\sqrt{\rho})c_{P}^{1/2}J^{-1/2}\right)$
$-(\lambda+\mu)g^{\alpha\beta}\Gamma^{\gamma}_{\alpha\beta}\frac{\mathrm{i}}{(c_{S}^{2}-c_{P}^{2})\sqrt{\rho}}\left((c_{P}^{2}-3c_{S}^{2})(\partial_{\gamma}\log\sqrt{\rho})c_{P}^{1/2}J^{-1/2}\right)$
$-c_{P}^{-1}\partial_{\alpha}c_{P}\mu g^{\alpha\beta}\frac{\mathrm{i}}{(c_{S}^{2}-c_{P}^{2})\sqrt{\rho}}\left((c_{P}^{2}-3c_{S}^{2})(\partial_{\beta}\log\sqrt{\rho})c_{P}^{1/2}J^{-1/2}\right)$
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.$
The above formula can be simplified to
$\begin{split}H_{1}=&-\mathrm{i}\sqrt{\rho}(c_{P}^{2}-3c_{S}^{2})c_{P}^{1/2}J^{-1/2}g^{\alpha\beta}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}\\ &+\mathrm{i}\left(c_{P}^{2}-3c_{S}^{2}-\frac{2(c_{P}^{2}-3c_{S}^{2})}{c_{P}^{2}-c_{S}^{2}}\right)\sqrt{\rho}c_{S}^{2}c_{P}^{1/2}J^{-1/2}g^{\alpha\beta}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}\log\sqrt{\rho}\\ &+\mathrm{i}g^{\alpha\beta}\Gamma^{\gamma}_{\alpha\beta}(c_{P}^{2}-3c_{S}^{2})c_{P}^{1/2}J^{-1/2}\partial_{\gamma}\log\sqrt{\rho}\\ &-\mathrm{i}\sqrt{\rho}(c_{P}^{2}-3c_{S}^{2})c_{P}^{-1/2}J^{-1/2}g^{\alpha\beta}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}c_{P}\\ &+\mathrm{i}\sqrt{\rho}\left(2c_{S}-\frac{4c_{S}(c_{P}^{2}-5c_{S}^{2})}{c_{P}^{2}-c_{S}^{2}}\right)c_{P}^{-1/2}J^{-1/2}g^{\alpha\beta}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}c_{S}\\ &+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.\end{split}$
Further, using the identities
$\partial^{2}_{33}\rho^{-1/2}=-\rho^{-1/2}\partial^{2}_{33}\log\sqrt{\rho}+\rho^{-1/2}(\partial_{3}\log\sqrt{\rho})^{2},$
$\partial_{\alpha}\partial_{\beta}\rho^{-1/2}=-\rho^{-1/2}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}+\rho^{-1/2}(\partial_{\alpha}\log\sqrt{\rho})(\partial_{\beta}\log\sqrt{\rho}),$
we have
$H_{2}=$ $(\lambda+2\mu)c_{P}^{-1}\left(2(\partial_{3}J^{-1/2})c_{P}^{-1/2}\partial_{3}\rho^{-1/2}+2J^{-1/2}(\partial_{3}c_{P}^{-1/2})\partial_{3}\rho^{-1/2}+J^{-1/2}c_{P}^{-1/2}\partial^{2}_{33}\rho^{-1/2}\right)$
$+\left(2(\lambda+2\mu)\partial_{3}\log\sqrt{\rho}+\partial_{3}(c_{P}^{2})\rho\right)c_{P}^{-1}J^{-1/2}c_{P}^{-1/2}(-\frac{1}{2}\rho^{-1/2}\partial_{3}\log\sqrt{\rho})$
$+2(\lambda+2\mu)\partial_{3}\log\sqrt{\rho}c_{P}^{-1}\partial_{3}(c_{P}^{-1/2}J^{-1/2})\rho^{-1/2}$
$-c_{P}^{-2}(\partial_{3}c_{P})(\lambda+2\mu)J^{-1/2}c_{P}^{-1/2}(-\rho^{-1/2}\partial_{3}\log\sqrt{\rho})$
$+g^{\alpha\beta}(\partial_{\beta}\mu)\left(\partial_{\alpha}J^{-1/2}c_{P}^{1/2}\rho^{-1/2}+J^{-1/2}\partial c_{P}^{1/2}\rho^{-1/2}+J^{-1/2}c_{P}^{1/2}\partial_{\alpha}\rho^{-1/2}\right)$
$+\mu g^{\alpha\beta}\left(2\partial_{\alpha}J^{-1/2}c_{P}^{1/2}\partial_{\beta}\rho^{-1/2}+2J^{-1/2}\partial_{\alpha}c_{P}^{1/2}\partial_{\beta}\rho^{-1/2}+J^{-1/2}c_{P}^{1/2}\partial_{\alpha}\partial_{\beta}\rho^{-1/2}\right)$
$+(c_{P}^{2}-2c_{S}^{2})\rho\partial_{3}\log\sqrt{\rho}c_{P}^{-1}g^{\alpha\beta}\frac{\partial g^{\alpha\beta}}{\partial\tau}c_{P}^{-1/2}J^{-1/2}\rho^{-1/2}$
$-4g^{\alpha\beta}\mu\partial_{\beta}c_{P}J^{-1/2}c_{P}^{-1/2}\partial_{\alpha}\log\sqrt{\rho}$
$-g^{\alpha\beta}(\partial_{\beta}c_{P})\mu J^{-1/2}c_{P}^{-1/2}(-\rho^{-1/2}\partial_{\alpha}\log\sqrt{\rho})$
$+\frac{1}{2}g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}(\lambda+2\mu)c_{P}^{-1}J^{-1/2}c_{P}^{-1/2}\left(-\rho^{-1/2}\partial_{3}\log\sqrt{\rho}\right)-g^{\gamma\delta}\Gamma^{\alpha}_{\gamma\delta}\mu c_{P}J^{-1/2}c_{P}^{-1/2}(-\rho^{-1/2}\partial_{\alpha}\log\sqrt{\rho})$
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}$
$=$ $2(\lambda+2\mu)c_{P}^{-1}(\partial_{3}J^{-1/2})c_{P}^{-1/2}(-\rho^{-1/2}\partial_{3}\log\sqrt{\rho})+(\lambda+2\mu)c_{P}^{-1}J^{-1/2}c_{P}^{-3/2}(\partial_{3}c_{P})(-\rho^{-1/2}\partial_{3}\log\sqrt{\rho})$
$+(\lambda+2\mu)c_{P}^{-1}J^{-1/2}c_{P}^{-1/2}\rho^{-1/2}(-\partial^{2}_{33}\log\sqrt{\rho}+(\partial_{3}\log\sqrt{\rho})^{2})$
$+\left(2(\lambda+2\mu)\partial_{3}\log\sqrt{\rho}+2c_{P}\partial_{3}c_{P}\rho\right)c_{P}^{-1}J^{-1/2}c_{P}^{-1/2}(-\frac{1}{2}\rho^{-1/2}\partial_{3}\log\sqrt{\rho})$
$+c_{P}^{-1/2}\sqrt{\rho}J^{-1/2}\partial_{3}\log\sqrt{\rho}\partial_{3}c_{P}+2c_{P}^{1/2}\sqrt{\rho}\partial_{3}J^{-1/2}\partial_{3}\log\sqrt{\rho}$
$+c_{P}^{-2}(\partial_{3}c_{P})(\lambda+2\mu)J^{-1/2}c_{P}^{-1/2}\rho^{-1/2}\partial_{3}\log\sqrt{\rho}$
$+2g^{\alpha\beta}c_{S}^{2}\rho\partial_{\beta}\log\sqrt{\rho}\left(\partial_{\alpha}J^{-1/2}c_{P}^{1/2}\rho^{-1/2}+\frac{1}{2}J^{-1/2}c_{P}^{-1/2}(\partial_{\alpha}c_{P})\rho^{-1/2}-J^{-1/2}c_{P}^{1/2}\rho^{-1/2}\partial_{\alpha}\log\sqrt{\rho}\right)$
$-2c_{s}(\partial_{\beta}c_{S})\rho^{1/2}J^{-1/2}g^{\alpha\beta}(\partial_{\alpha}\log\sqrt{\rho})$
$+\mu g^{\alpha\beta}\Big{(}-2\partial_{\alpha}J^{-1/2}c_{P}^{1/2}\rho^{-1/2}\partial_{\beta}\log\sqrt{\rho}+J^{-1/2}\rho^{-1/2}c_{P}^{-1/2}\partial_{\alpha}c_{P}\partial_{\beta}\log\sqrt{\rho}$
$\quad\quad+\rho^{-1/2}J^{-1/2}c_{P}^{1/2}(-\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}+(\partial_{\alpha}\log\sqrt{\rho})(\partial_{\beta}\log\sqrt{\rho}))\Big{)}$
$+(c_{P}^{2}-2c_{S}^{2})\rho\partial_{3}\log\sqrt{\rho}c_{P}^{-1}g^{\alpha\beta}\frac{\partial g^{\alpha\beta}}{\partial\tau}c_{P}^{-1/2}J^{-1/2}\rho^{-1/2}$
$-3g^{\alpha\beta}\mu\partial_{\beta}c_{P}J^{-1/2}c_{P}^{-1/2}\partial_{\alpha}\log\sqrt{\rho}$
$-\frac{1}{2}g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}(\lambda+2\mu)c_{P}^{-1}J^{-1/2}c_{P}^{-1/2}\rho^{-1/2}\partial_{3}\log\sqrt{\rho}+g^{\gamma\delta}\Gamma^{\alpha}_{\gamma\delta}\mu c_{P}J^{-1/2}c_{P}^{-1/2}\rho^{-1/2}\partial_{\alpha}\log\sqrt{\rho}$
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.$
The above formula simplifies to
$H_{2}=$ $-J^{-1/2}c_{P}^{1/2}\rho^{1/2}\partial^{2}_{33}\log\sqrt{\rho}-c_{S}^{2}\rho^{1/2}J^{-1/2}c_{P}^{1/2}g^{\alpha\beta}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}$
$-c_{S}^{2}\rho^{1/2}J^{-1/2}c_{P}^{1/2}g^{\alpha\beta}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}\log\sqrt{\rho}-c_{P}c_{P}^{-1/2}\rho^{1/2}J^{-1/2}(\partial_{3}\log\sqrt{\rho})^{2}$
$+g^{\gamma\delta}\Gamma^{\alpha}_{\gamma\delta}c_{S}^{2}c_{P}^{1/2}J^{-1/2}\rho^{1/2}\partial_{\alpha}\log\sqrt{\rho}+c_{P}^{-1/2}c_{P}^{-2}(3c_{P}^{2})\rho^{1/2}J^{-1/2}\partial_{3}c_{P}\partial_{3}\log\sqrt{\rho}$
$-3c_{S}^{2}c_{P}^{-1/2}\rho J^{-1/2}g^{\alpha\beta}\partial_{\alpha}c_{P}\partial_{\beta}\log\sqrt{\rho}-2c_{s}(\partial_{\beta}c_{S})\rho^{1/2}J^{-1/2}g^{\alpha\beta}(\partial_{\alpha}\log\sqrt{\rho})$
$+\frac{1}{2}g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}(c_{P}^{2}-4c_{S}^{2})c_{P}^{-1}J^{-1/2}c_{P}^{-1/2}\rho^{1/2}\partial_{3}\log\sqrt{\rho}$
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.$
Then one can write
$-\mathrm{i}H_{1}-H_{2}=$ $-\sqrt{\rho}(c_{P}^{2}-4c_{S}^{2})c_{P}^{1/2}J^{-1/2}g^{\alpha\beta}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}$
$+\left(c_{P}^{2}-2c_{S}^{2}-\frac{2c_{S}^{2}(c_{P}^{2}-3c_{S}^{2})}{c_{P}^{2}-c_{S}^{2}}\right)\sqrt{\rho}c_{P}^{1/2}J^{-1/2}g^{\alpha\beta}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}\log\sqrt{\rho}$
$+g^{\alpha\beta}\Gamma^{\gamma}_{\alpha\beta}(c_{P}^{2}-4c_{S}^{2})c_{P}^{1/2}J^{-1/2}\partial_{\gamma}\log\sqrt{\rho}+J^{-1/2}c_{P}^{1/2}\rho^{1/2}\partial^{2}_{33}\log\sqrt{\rho}$
$-\frac{1}{2}g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}(c_{P}^{2}-4c_{S}^{2})\rho^{1/2}c_{P}^{-1}J^{-1/2}c_{P}^{-1/2}\partial_{3}\log\sqrt{\rho}$
$+c_{P}^{-1/2}(1+c_{P}^{-2}c_{S}^{2})\rho^{1/2}J^{-1/2}\partial_{3}c_{P}\partial_{3}\log\sqrt{\rho}$
$+\sqrt{\rho}\left(\frac{8c_{S}^{3}}{c_{P}^{2}-c_{S}^{2}}\right)c_{P}^{1/2}J^{-1/2}g^{\alpha\beta}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}c_{S}-c_{P}^{2}c_{P}^{-1/2}\rho^{1/2}J^{-1/2}g^{\alpha\beta}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}c_{P}$
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.$
Using the formulas
(11) $\nabla_{3}\nabla_{3}\log\sqrt{\rho}=\partial^{2}_{33}\log\sqrt{\rho}-c_{P}^{-1}\partial_{3}c_{P}\partial_{3}\log\sqrt{\rho}+c_{P}g^{\alpha\beta}\partial_{\beta}c_{P}\partial_{\alpha}\log\sqrt{\rho},$
(12) $\nabla_{\alpha}\nabla_{\beta}\log\sqrt{\rho}=\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}+\frac{1}{2}c_{P}^{-2}\frac{g_{\alpha\beta}}{\partial\tau}\partial_{3}\log\sqrt{\rho}-\Gamma^{\gamma}_{\alpha\beta}\partial_{\gamma}\log\sqrt{\rho},$
we obtain
$-\mathrm{i}H_{1}-H_{2}$
$=$ $\rho^{1/2}J^{-1/2}c_{P}^{1/2}\Big{(}-(c_{P}^{2}-4c_{S}^{2})g^{\alpha\beta}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}+(c_{P}^{2}-4c_{S}^{2})g^{\alpha\beta}\Gamma^{\gamma}_{\alpha\beta}\partial_{\gamma}\log\sqrt{\rho}$
$-(c_{P}^{2}-4c_{S}^{2})\frac{1}{2}c_{P}^{-2}\frac{\partial g_{\alpha\beta}}{\partial\tau}g^{\alpha\beta}\partial_{3}\log\sqrt{\rho}-(c_{P}^{2}-4c_{S}^{2})c_{P}^{-2}\partial^{2}_{33}\log\sqrt{\rho}+(c_{P}^{2}-4c_{S}^{2})c_{P}^{-2}\partial_{3}c_{P}\partial_{3}\log\sqrt{\rho}$
$-(c_{P}^{2}-4c_{S}^{2})c_{P}^{-1}g^{\alpha\beta}\partial_{\beta}c_{P}\partial_{\alpha}\log\sqrt{\rho}\Bigg{)}$
$+(2c_{P}^{2}-4c_{S}^{2})\rho^{1/2}J^{-1/2}c_{P}^{1/2}\left(c_{P}^{-2}\partial^{2}_{33}\log\sqrt{\rho}-c_{P}^{-3}\partial_{3}c_{P}\partial_{3}\log\sqrt{\rho}+c_{P}^{-1}g^{\alpha\beta}\partial_{\beta}c_{P}\partial_{\alpha}\log\sqrt{\rho}\right)$
$+c_{P}c_{P}^{-1/2}\rho^{1/2}J^{-1/2}(\partial_{3}\log\sqrt{\rho})^{2}$
$+\sqrt{\rho}c_{P}^{1/2}J^{-1/2}\left(c_{P}^{2}-4c_{S}^{2}+\frac{4c_{S}^{4}}{c_{P}^{2}-c_{S}^{2}}\right)g^{\alpha\beta}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}\log\sqrt{\rho}$
$+\sqrt{\rho}c_{P}^{-3/2}J^{-1/2}\partial_{3}c_{P}^{2}\partial_{3}\log\sqrt{\rho}$
$-c_{P}^{1/2}\rho^{1/2}J^{-1/2}g^{\alpha\beta}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}c_{P}^{2}+\sqrt{\rho}\left(\frac{16c_{S}^{3}}{c_{P}^{2}-c_{S}^{2}}\right)c_{P}^{1/2}J^{-1/2}g^{\alpha\beta}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}c_{S}$
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.$
$=$ $\rho^{1/2}J^{-1/2}c_{P}^{1/2}\left(-(c_{P}^{2}-4c_{S}^{2})\Delta\log\sqrt{\rho}\right)+(2c_{P}^{2}-4c_{S}^{2})\rho^{1/2}J^{-1/2}c_{P}^{1/2}c_{P}^{-2}\nabla^{2}_{33}\log\sqrt{\rho}$
$+\rho^{1/2}J^{-1/2}c_{P}^{1/2}\left(c_{P}^{2}-4c_{S}^{2}+\frac{4c_{S}^{4}}{c_{P}^{2}-c_{S}^{2}}\right)|\nabla\log\sqrt{\rho}|^{2}+4\rho^{1/2}J^{-1/2}c_{P}^{1/2}\frac{c_{S}^{2}}{c_{P}^{2}}\frac{c_{P}^{2}-2c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}|\partial_{3}\log\sqrt{\rho}|^{2}$
$+\rho^{1/2}J^{-1/2}c_{P}^{1/2}\nabla\log\sqrt{\rho}\cdot\left(-\nabla c_{P}^{2}+\frac{8c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}\nabla c_{S}^{2}\right)$
$+\rho^{1/2}J^{-1/2}c_{P}^{-3/2}\partial_{3}\log\sqrt{\rho}\left(2\partial_{3}c_{P}^{2}-\frac{8c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}\partial_{3}c_{S}^{2}\right)$
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.$
Inserting $a_{13}:=c_{P}^{1/2}\rho^{-1/2}J^{-1/2}B$ into the equation (10), we get an equation for $B$ along the unit-speed geodesic $\gamma(s)$ in $(\Omega,g_{P})$,
$2\mathrm{i}\frac{\mathrm{d}B}{\mathrm{d}s}=M_{ij}\dot{\gamma}^{i}\dot{\gamma}^{j}+C,$
where $M$ is a second order tensor
$M=$ $-(c_{P}^{2}-4c_{S}^{2})\Delta\log\sqrt{\rho}g_{P}+\frac{2c_{P}^{2}-4c_{S}^{2}}{c_{P}^{2}}\nabla^{2}\log\sqrt{\rho}$
$+\left(c_{P}^{2}-4c_{S}^{2}+\frac{4c_{S}^{4}}{c_{P}^{2}-c_{S}^{2}}\right)|\nabla\log\sqrt{\rho}|^{2}g_{P}+\frac{c_{S}^{2}}{c_{P}^{2}}\frac{c_{P}^{2}-2c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}\nabla\log\sqrt{\rho}\otimes\nabla\log\sqrt{\rho}$
$+\nabla\log\sqrt{\rho}\cdot\left(-\nabla c_{P}^{2}+\frac{8c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}\nabla c_{S}^{2}\right)g_{P}+\frac{1}{c_{P}^{2}}\nabla\log\sqrt{\rho}\otimes\left(2\nabla c_{P}^{2}-\frac{8c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}\nabla c_{S}^{2}\right),$
and $C$ does not depend on $\rho$. Recall that $g_{P}=c_{P}^{-2}g_{E}$ is the metric induced by the $P$-wavespeed.
If we solve $B$ along a geodesic $\gamma$ connecting two points $a\in\partial\Omega$ to $b\in\partial\Omega$, then
$B(b)-B(a)=\frac{1}{2\mathrm{i}}\int_{\gamma}(M_{ij}(\gamma(s))\dot{\gamma}^{i}(s)\dot{\gamma}^{j}(s)+C)\mathrm{d}s.$
By carefully analyzing the reflection of waves (with possible mode conversions) (see, for example, [28, 8]) at the boundary or using the argument in [20], one can show that the Dirichlet-to-Neumann map $\Lambda_{\rho}$ determines $B$ on $\partial\Omega$. Similar case is analyzed for acoustic waves in [29]. Therefore we can recover the longitudinal geodesic ray transform of $M$
$\int_{\gamma}M_{ij}(\gamma(s))\dot{\gamma}^{i}(s)\dot{\gamma}^{j}(s)\mathrm{d}s.$
Here $\gamma$ is an arbitrary geodesic in $(\Omega,g_{P})$ connecting two boundary points.
We remark here that the same result was obtained in [20, 6], and then the uniqueness of $\rho$ can be proved assuming, for example, $\lambda=2\mu$ only at isolated points. Here one needs to use the injectivity, up to natural obstructions, of the longitudinal geodesic ray transform (cf., for example, [21, 15, 27]).
4. S-waves and proof of the main theorem
In this section, we will use solutions of the form (4) that represent $S$-waves to derive new results. We work under the ray coordinates associated with the metric $g_{S}=c_{S}^{-2}\mathrm{d}s^{2}$. Now we take
$|\nabla\varphi|^{2}=c_{S}^{-2}$
and $\langle\mathbf{a}_{0},\nabla\varphi\rangle=0$ (or equivalently $a_{03}=0$). Then the identity (7) is satisfied.
We calculate
$\partial_{\alpha}(\lambda a_{0k}\varphi_{;\ell}g^{k\ell})+\lambda a_{0k}\varphi_{;\ell}g^{k\ell}\partial_{m}g_{\alpha j}g^{jm}+\partial_{m}(\mu a_{0\alpha}\varphi_{;j}+\mu a_{0j}\varphi_{;\alpha})g^{jm}$
$=$ $\partial_{3}(\mu a_{0\alpha}\varphi_{;3})g^{33}$
$=$ $c_{S}^{-2}\partial_{3}(\mu a_{0\alpha}),$
$\lambda a_{0k;\ell}g^{k\ell}\varphi_{;\alpha}+\mu(a_{0\alpha;j}+a_{0j;\alpha})\varphi_{;m}g^{jm}$
$=$ $c_{S}^{-2}\mu(\partial_{3}a_{0\alpha}-2\Gamma^{\beta}_{3\alpha}a_{0\beta})$
$=$ $c_{S}^{-2}\mu\partial_{3}a_{0\alpha}-c_{S}^{-2}\mu g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}a_{0\beta}$
$\Gamma^{n}_{\alpha m}(\lambda a_{0k}\varphi_{;\ell}g^{k\ell}g_{nj}+\mu a_{0n}\varphi_{;j}+\mu a_{0j}\varphi_{;n})g^{jm}$
$=$ $g^{33}\Gamma^{\gamma}_{\alpha 3}\mu a_{0\gamma}+g^{\beta\gamma}\Gamma^{3}_{\alpha\beta}\mu a_{0\gamma}=0,$
$\Gamma^{n}_{jm}(\lambda a_{0k}\varphi_{;\ell}g^{k\ell}g_{n\alpha}+\mu a_{0n}\varphi_{;\alpha}+\mu a_{0\alpha}\varphi_{;n})g^{jm}$
$=$ $(g^{33}\Gamma^{3}_{33}+g^{\beta\gamma}\Gamma^{3}_{\beta\gamma})\mu a_{0\alpha}$
$=$ $c_{S}^{-3}\partial_{3}c_{S}\mu a_{0\alpha}-\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}c_{S}^{-2}\mu a_{0\alpha}.$
As for the $P$-waves, we can also take $\mathbf{a}_{j}$ idependent of $t$. Equating the $\mathcal{O}(\varrho)$-terms in the equation (6) with $i=\alpha$, we obtain
$c_{S}^{-2}\left(2\mu\partial_{3}a_{0\alpha}+(\partial_{3}\mu)a_{0\alpha}-\mu g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}a_{0\beta}+\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}c_{S}^{-2}\mu a_{0\alpha}-c_{S}^{-1}\partial_{3}c_{S}\mu a_{0\alpha}\right)=0.$
Introducing the vector $\eta$ such that $\eta_{3}=0$ and
(13) $\frac{\partial\eta_{\alpha}}{\partial\tau}-\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\alpha\beta}}{\partial\tau}\eta_{\gamma}+c_{S}^{-1}\partial_{3}c_{S}\eta_{\alpha}=0.$
and denote
$u_{\alpha}=A_{S}c_{S}\eta_{\alpha}.$
If we consider the equation (13) as an ODE along a geodesic $\gamma$ with respect to the metric $g_{S}$, we can see that $\eta$ is a parallel vector field along $\gamma$. So one can take $|\eta|^{2}_{g_{S}}=1$. We obtain the following equation for $A_{S}$,
$c_{S}^{-2}\left(2\mu c_{S}\partial_{3}A_{S}+(\partial_{3}\mu)c_{S}A_{S}-\partial_{3}c_{S}\mu A_{S}+\frac{1}{2}\mu g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}c_{S}A_{S}\right)=0.$
Similar to the $P$-waves, the solution is
$A_{S}=cJ^{-1/2}c_{S}^{-1/2}\rho^{-1/2}$
with some constant $c$. As in previous section, we take $c=1$. Observe that one can not get any information about $\rho$ in the interior of $\Omega$ from the leading order term of the amplitude at the boundary, which is also pointed out in [20].
Next we determine $a_{13}$. For this, we also calculate
$\partial_{3}(\lambda a_{0k}\varphi_{;\ell}g^{k\ell})+\lambda a_{0k}\varphi_{;\ell}g^{k\ell}\partial_{m}g_{3j}g^{jm}+\partial_{m}(\mu a_{03}\varphi_{;j}+\mu a_{0j}\varphi_{;3})g^{jm}$
$=$ $g^{\alpha\beta}\partial_{\beta}(\mu a_{0\alpha}),$
$\lambda a_{0k;\ell}g^{k\ell}\varphi_{;3}+\mu(a_{03;j}+a_{0j;3})\varphi_{;m}g^{jm}$
$=$ $\lambda(\partial_{\beta}a_{0\alpha}-\Gamma^{\gamma}_{\alpha\beta}a_{0\gamma})g^{\alpha\beta}-\lambda\Gamma^{\alpha}_{33}a_{0\alpha}g^{33}+2g^{33}\mu(-\Gamma_{33}^{\alpha}a_{0\alpha})$
$=$ $\lambda g^{\alpha\beta}\partial_{\beta}a_{0\alpha}-g^{\alpha\beta}\lambda\Gamma^{\gamma}_{\alpha\beta}a_{0\gamma}+(\lambda+2\mu)c_{S}^{-1}g^{\alpha\beta}\partial_{\beta}c_{S}a_{0\alpha},$
$\Gamma^{n}_{3m}(\lambda a_{0k}\varphi_{;\ell}g^{k\ell}g_{nj}+\mu a_{0n}\varphi_{;j}+\mu a_{0j}\varphi_{;n})g^{jm}$
$=$ $g^{33}\Gamma^{\alpha}_{33}\mu a_{0\alpha}+g^{\alpha\beta}\Gamma^{3}_{3\beta}\mu a_{0\alpha}$
$=$ $-c_{S}^{-1}g^{\alpha\beta}\partial_{\beta}c_{S}\mu a_{0\alpha}+g^{\alpha\beta}\mu c_{S}^{-1}\partial_{\beta}c_{S}a_{0\alpha}=0,$
$\Gamma^{n}_{jm}(\lambda a_{0k}\varphi_{;\ell}g^{k\ell}g_{n3}+\mu a_{0n}\varphi_{;3}+\mu a_{03}\varphi_{;n})g^{jm}$
$=$ $g^{jm}\Gamma_{jm}^{\alpha}\mu a_{0\alpha}$
$=$ $-c_{S}^{-1}\mu g^{\alpha\beta}\partial_{\beta}c_{S}a_{0\alpha}+g^{\beta\gamma}\Gamma^{\alpha}_{\beta\gamma}\mu a_{0\alpha}.$
Equating the $\mathcal{O}(\varrho)$-terms in the equation (6) with $i=3$, we obtain
$\begin{split}\rho a_{13}-(\lambda+2\mu)c_{S}^{-2}a_{13}\\ +\mathrm{i}\Big{(}g^{\alpha\beta}\partial_{\beta}(\mu a_{0\alpha})+\lambda g^{\alpha\beta}\partial_{\beta}a_{0\alpha}-\lambda g^{\alpha\beta}\Gamma^{\gamma}_{\alpha\beta}a_{0\gamma}+(\lambda+2\mu)c_{S}^{-1}g^{\alpha\beta}\partial_{\beta}c_{S}a_{0\alpha}\\ +c_{S}^{-1}\mu g^{\alpha\beta}\partial_{\beta}c_{S}a_{0\alpha}-g^{\beta\gamma}\Gamma^{\alpha}_{\beta\gamma}\mu a_{0\alpha}\Big{)}=0,\end{split}$
which simplifies to
$\begin{split}\rho a_{13}-(\lambda+2\mu)c_{S}^{-2}a_{13}\\ +\mathrm{i}\Big{(}(\lambda+\mu)g^{\alpha\beta}\partial_{\beta}a_{0\alpha}+g^{\alpha\beta}(\partial_{\beta}\mu)a_{0\alpha}+(\lambda+3\mu)c_{S}^{-1}g^{\alpha\beta}\partial_{\beta}c_{S}a_{0\alpha}-(\lambda+\mu)g^{\beta\gamma}\Gamma^{\alpha}_{\beta\gamma}a_{0\alpha}\Big{)}=0.\end{split}$
So
$\begin{split}a_{13}=&\frac{\mathrm{i}}{c_{P}^{2}c_{S}^{-2}-1}\left(\frac{\lambda+\mu}{\rho}g^{\alpha\beta}\partial_{\beta}a_{0\alpha}+g^{\alpha\beta}\frac{\partial_{\beta}\mu}{\rho}a_{0\alpha}+\frac{\lambda+3\mu}{\rho}c_{S}^{-1}g^{\alpha\beta}\partial_{\beta}c_{S}a_{0\alpha}-\frac{\lambda+\mu}{\rho}g^{\beta\gamma}\Gamma^{\alpha}_{\beta\gamma}a_{0\alpha}\right)\\ =&\frac{\mathrm{i}}{c_{P}^{2}c_{S}^{-2}-1}\Bigg{(}(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\partial_{\beta}a_{0\gamma}+g^{\alpha\beta}\partial_{\beta}c_{S}^{2}a_{0\gamma}+2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}+(c_{P}^{2}+c_{S}^{2})c_{S}^{-1}g^{\beta\gamma}\partial_{\beta}c_{S}a_{0\gamma}\\ &\quad\quad\quad\quad-(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\Gamma^{\alpha}_{\beta\gamma}a_{0\alpha}\Bigg{)}.\end{split}$
To continue, we calculate
$\partial_{\alpha}a_{13}=$ $\frac{\mathrm{i}}{c_{P}^{2}c_{S}^{-2}-1}\Bigg{(}(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\partial_{\alpha}\partial_{\beta}a_{0\gamma}+\partial_{\alpha}(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\partial_{\beta}a_{0\gamma}+(c_{P}^{2}-c_{S}^{2})\partial_{\alpha}g^{\beta\gamma}\partial_{\beta}a_{0\gamma}$
$\quad\quad+g^{\beta\gamma}(\partial_{\beta}c_{S}^{2})\partial_{\alpha}a_{0\gamma}+2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\partial_{\alpha}a_{0\gamma}+2\partial_{\alpha}g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$\quad\quad+2g^{\beta\gamma}\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}+2g^{\beta\gamma}c_{S}^{2}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$\quad\quad+(c_{P}^{2}+c_{S}^{2})c_{S}^{-1}g^{\beta\gamma}\partial_{\beta}c_{S}\partial_{\alpha}a_{0\gamma}-(c_{P}^{2}+c_{S}^{2})g^{\beta\gamma}\Gamma^{\delta}_{\beta\gamma}\partial_{\alpha}a_{0\delta}$
$\quad\quad-\frac{1}{c_{S}^{2}}\frac{c_{S}^{2}\partial_{\alpha}c_{P}^{2}+c_{P}^{2}\partial_{\alpha}c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\partial_{\beta}a_{0\gamma}-\frac{2}{c_{S}^{2}}\frac{c_{S}^{2}\partial_{\alpha}c_{P}^{2}+c_{P}^{2}\partial_{\alpha}c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}\Bigg{)}$
$+\frac{1}{\sqrt{\rho}}\times\text{terms that do not depend on $\rho$}$
$=$ $\frac{\mathrm{i}}{c_{P}^{2}c_{S}^{-2}-1}\Bigg{(}(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\partial_{\alpha}\partial_{\beta}a_{0\gamma}-\frac{c_{P}^{2}+c_{S}^{2}}{c_{S}^{2}}\partial_{\alpha}(c_{S}^{2})g^{\beta\gamma}\partial_{\beta}a_{0\gamma}+(c_{P}^{2}-c_{S}^{2})\partial_{\alpha}g^{\beta\gamma}\partial_{\beta}a_{0\gamma}$
$\quad\quad+2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\partial_{\alpha}a_{0\gamma}+2\partial_{\alpha}g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$\quad\quad+2g^{\beta\gamma}c_{S}^{2}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$\quad\quad+(c_{P}^{2}+3c_{S}^{2})c_{S}^{-1}g^{\beta\gamma}\partial_{\beta}c_{S}\partial_{\alpha}a_{0\gamma}-(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\Gamma^{\delta}_{\beta\gamma}\partial_{\alpha}a_{0\delta}$
$\quad\quad+2g^{\beta\gamma}\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}-2\frac{c_{S}^{2}\partial_{\alpha}c_{P}^{2}+c_{P}^{2}\partial_{\alpha}c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}\Bigg{)}$
$+\frac{1}{\sqrt{\rho}}\times\text{terms that do not depend on $\rho$}.$
We proceed with the tedious calculation
$\partial_{\alpha}(\lambda a_{1k}\varphi_{;\ell}g^{k\ell})+\lambda a_{1k}\varphi_{;\ell}g^{k\ell}\partial_{m}g_{\alpha j}g^{jm}+\partial_{m}(\mu a_{1\alpha}\varphi_{;j}+\mu a_{1j}\varphi_{;\alpha})g^{jm}$
$=$ $\partial_{\alpha}(\lambda a_{13}g^{33})+\lambda a_{13}g^{33}g^{\beta\gamma}\partial_{\beta}g_{\alpha\gamma}+\partial_{3}(\mu a_{1\alpha})g^{33}$
$=$ $\partial_{\alpha}(\lambda a_{13}c_{S}^{-2})+\lambda a_{13}c_{S}^{-2}g^{\beta\gamma}\partial_{\beta}g_{\alpha\gamma}+\partial_{3}(\mu a_{1\alpha})c_{S}^{-2},$
$\lambda a_{1k;\ell}g^{k\ell}\varphi_{;\alpha}+\mu(a_{1\alpha;j}+a_{1j;\alpha})\varphi_{;m}g^{jm}$
$=$ $\mu(\partial_{3}a_{1\alpha})g^{33}+\mu(\partial_{\alpha}a_{13})g^{33}-2\mu\Gamma^{i}_{\alpha 3}a_{1i}g^{33}$
$=$ $\mu c_{S}^{-2}\partial_{3}a_{1\alpha}+\mu c_{S}^{-2}\partial_{\alpha}a_{13}-2\mu c_{S}^{-2}(\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}a_{1\beta}+c_{S}^{-1}\partial_{\alpha}c_{S}a_{13})$
$=$ $\mu c_{S}^{-2}\partial_{3}a_{1\alpha}-\mu c_{S}^{-2}g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}a_{1\beta}+\mu c_{S}^{-2}\partial_{\alpha}a_{13}-2\mu c_{S}^{-3}\partial_{\alpha}c_{S}a_{13},$
$\Gamma^{n}_{\alpha m}(\lambda a_{1k}\varphi_{;\ell}g^{k\ell}g_{nj}+\mu a_{1n}\varphi_{;j}+\mu a_{1j}\varphi_{;n})g^{jm}$
$=$ $\Gamma_{\alpha n}^{n}\lambda a_{13}g^{33}+\Gamma^{n}_{\alpha 3}\mu a_{1n}g^{33}+\Gamma^{3}_{\alpha m}\mu a_{1j}g^{jm}$
$=$ $(\Gamma_{\alpha 3}^{3}+\Gamma_{\alpha\beta}^{\beta})\lambda a_{13}g^{33}+\Gamma^{3}_{\alpha 3}\mu a_{13}g^{33}+\Gamma^{\beta}_{\alpha 3}\mu a_{1\beta}g^{33}+\Gamma^{3}_{\alpha 3}\mu a_{13}g^{33}+\Gamma^{3}_{\alpha\beta}\mu a_{1\gamma}g^{\beta\gamma}$
$=$ $(c_{S}^{-1}\partial_{\alpha}c_{S}+\Gamma^{\beta}_{\alpha\beta})\lambda c_{S}^{-2}a_{13}+2\mu c_{S}^{-3}\partial_{\alpha}c_{S}a_{13},$
$\Gamma^{n}_{jm}(\lambda a_{1k}\varphi_{;\ell}g^{k\ell}g_{n\alpha}+\mu a_{1n}\varphi_{;\alpha}+\mu a_{1\alpha}\varphi_{;n})g^{jm}$
$=$ $g^{jm}\Gamma^{\beta}_{jm}\lambda a_{13}g^{33}g_{\alpha\beta}+(g^{33}\Gamma^{3}_{33}+g^{\beta\gamma}\Gamma^{3}_{\beta\gamma})\mu a_{1\alpha}$
$=$ $-\lambda c_{S}^{-3}\partial_{\alpha}c_{S}a_{13}+c_{S}^{-2}g^{\gamma\delta}g_{\alpha\beta}\Gamma^{\beta}_{\gamma\delta}\lambda a_{13}+c_{S}^{-3}\partial_{3}c_{S}\mu a_{1\alpha}-\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}c_{S}^{-2}\mu a_{1\alpha},$
and
$\partial_{m}(\lambda a_{0k;\ell}g^{k\ell}g_{\alpha j}+\mu a_{0\alpha;j}+\mu a_{0j;\alpha})g^{jm}$
$=$ $g^{\beta\gamma}\partial_{\beta}(\lambda g^{33}(-\Gamma_{33}^{\delta}a_{0\delta})g_{\alpha\gamma}+\lambda(\partial_{\varepsilon}a_{0\delta}-\Gamma_{\varepsilon\delta}^{\eta}a_{0\eta})g^{\delta\varepsilon}g_{\alpha\gamma})$
$+g^{jm}\partial_{m}(\mu(\partial_{j}a_{0\alpha}-\Gamma^{\beta}_{\alpha j}a_{0\beta}))+g^{jm}\partial_{m}(\mu(\partial_{\alpha}a_{0j}-\Gamma^{\beta}_{\alpha j}a_{0\beta}))$
$=$ $g^{\beta\gamma}\partial_{\beta}(\lambda(\partial_{\varepsilon}a_{0\delta})g^{\delta\varepsilon}g_{\alpha\gamma})-g^{\beta\gamma}\partial_{\beta}(\lambda g^{33}\Gamma^{\delta}_{33}a_{0\delta}g_{\alpha\gamma}+\lambda\Gamma^{\eta}_{\varepsilon\delta}a_{0\eta}g^{\delta\varepsilon}g_{\alpha\gamma})$
$+g^{33}\partial_{3}(\mu\partial_{3}a_{0\alpha})+g^{\beta\gamma}\partial_{\beta}(\mu\partial_{\gamma}a_{0\alpha})+g^{\beta\gamma}\partial_{\gamma}(\mu\partial_{\alpha}a_{0\beta})-2g^{33}\partial_{3}(\mu\Gamma^{\beta}_{\alpha 3}a_{0\beta})-2g^{\delta\gamma}\partial_{\delta}(\mu\Gamma^{\beta}_{\alpha\gamma}a_{0\beta})$
$=$ $\partial_{\alpha}\lambda g^{\delta\varepsilon}\partial_{\varepsilon}a_{0\delta}+\lambda g^{\delta\varepsilon}\partial_{\alpha}\partial_{\varepsilon}a_{0\delta}+\lambda\partial_{\varepsilon}a_{0\delta}\partial_{\alpha}g^{\delta\varepsilon}+\lambda g^{\beta\gamma}g^{\delta\varepsilon}\partial_{\beta}g_{\alpha\gamma}\partial_{\varepsilon}a_{0\delta}$
$+(\partial_{\alpha}\lambda)c_{S}^{-1}g^{\delta\beta}\partial_{\beta}c_{S}a_{0\delta}+\lambda c_{S}^{-1}g^{\delta\beta}\partial_{\beta}c_{S}\partial_{\alpha}a_{0\delta}-\partial_{\alpha}\lambda g^{\delta\varepsilon}\Gamma^{\eta}_{\varepsilon\delta}a_{0\eta}-\lambda g^{\delta\varepsilon}\Gamma^{\eta}_{\varepsilon\delta}\partial_{\alpha}a_{0\eta}$
$+c_{S}^{-2}\partial_{3}\mu\partial_{3}a_{0\alpha}+c_{S}^{-2}\mu\partial^{2}_{33}a_{0\alpha}+g^{\beta\gamma}\partial_{\beta}\mu\partial_{\gamma}a_{0\alpha}+\mu g^{\beta\gamma}\partial_{\beta}\partial_{\gamma}a_{0\alpha}+g^{\beta\gamma}\partial_{\gamma}\mu\partial_{\alpha}a_{0\beta}+\mu g^{\beta\gamma}\partial_{\alpha}\partial_{\gamma}a_{0\beta}$
$-2c_{S}^{-2}\mu\Gamma^{\beta}_{\alpha 3}\partial_{3}a_{0\beta}-2g^{\delta\gamma}\mu\Gamma^{\beta}_{\alpha\gamma}\partial_{\delta}a_{0\beta}-2c_{S}^{-2}\partial_{3}\mu\Gamma^{\beta}_{\alpha 3}a_{0\beta}-2g^{\delta\gamma}\partial_{\delta}\mu\Gamma^{\beta}_{\alpha\gamma}a_{0\beta},$
$\Gamma_{\alpha m}^{n}(\lambda a_{0k;\ell}g^{k\ell}g_{nj}+\mu a_{0n;j}+\mu a_{0j;n})g^{jm}$
$=$ $\Gamma^{n}_{\alpha n}\lambda g^{\beta\gamma}\partial_{\gamma}a_{0\beta}+g^{jm}\Gamma^{\beta}_{\alpha m}\mu\partial_{j}a_{0\beta}+g^{\beta\gamma}\mu\Gamma^{n}_{\alpha\gamma}\partial_{n}a_{0\beta}+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}$
$=$ $(\Gamma^{3}_{3\alpha}+\Gamma^{\delta}_{\alpha\delta})\lambda g^{\beta\gamma}\partial_{\gamma}a_{0\beta}+g^{33}\Gamma^{\beta}_{\alpha 3}\mu\partial_{3}a_{0\beta}+g^{\gamma\delta}\Gamma^{\beta}_{\alpha\gamma}\mu\partial_{\delta}a_{0\beta}+g^{\beta\gamma}\mu\Gamma^{3}_{\alpha\gamma}\partial_{3}a_{0\beta}+g^{\beta\gamma}\Gamma^{\delta}_{\alpha\gamma}\mu\partial_{\delta}a_{0\beta}$
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}$
$=$ $(\Gamma^{3}_{3\alpha}+\Gamma^{\delta}_{\alpha\delta})\lambda g^{\beta\gamma}\partial_{\gamma}a_{0\beta}+g^{\gamma\delta}\Gamma^{\beta}_{\alpha\gamma}\mu\partial_{\delta}a_{0\beta}+g^{\beta\gamma}\Gamma^{\delta}_{\alpha\gamma}\mu\partial_{\delta}a_{0\beta}+\sqrt{\rho}\times\text{terms that do notdepend on $\rho$},$
$\Gamma_{jm}^{n}(\lambda a_{0k;\ell}g^{k\ell}g_{n\alpha}+\mu a_{0n;\alpha}+\mu a_{0\alpha;n})g^{jm}$
$=$ $g^{jm}\Gamma^{\beta}_{jm}g_{\alpha\beta}\lambda\partial_{\delta}a_{0\gamma}g^{\gamma\delta}+g^{jm}\Gamma^{\beta}_{jm}\mu\partial_{\alpha}a_{0\beta}+g^{jm}\Gamma^{\beta}_{jm}\mu\partial_{\beta}a_{0\alpha}+g^{jm}\Gamma^{3}_{jm}\mu\partial_{3}a_{0\alpha}$
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}$
$=$ ${g^{33}\Gamma^{\beta}_{33}g_{\alpha\beta}\lambda\partial_{\delta}a_{0\gamma}g^{\gamma\delta}}+{g^{\eta\varepsilon}\Gamma^{\beta}_{\eta\varepsilon}g_{\alpha\beta}\lambda\partial_{\delta}a_{0\gamma}g^{\gamma\delta}}+\mu(g^{33}\Gamma^{\beta}_{33}+g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta})(\partial_{\alpha}a_{0\beta}+\partial_{\beta}a_{0\alpha})$
$+(g^{33}\Gamma^{3}_{33}+g^{\beta\gamma}\Gamma^{3}_{\beta\gamma})\mu\partial_{3}a_{0\alpha}+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.$
Collecting $\mathcal{O}(1)$ terms in equation (6) with $i=\alpha$, we obtain
(14) $\mathrm{i}c_{S}^{-2}\left(2\mu\partial_{3}a_{1\alpha}+(\partial_{3}\mu)a_{1\alpha}-\mu g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}a_{1\beta}+\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}c_{S}^{-2}\mu a_{1\alpha}-c_{S}^{-1}\partial_{3}c_{S}\mu a_{1\alpha}\right)=\rho H,$
where
$H=$ ${(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\partial_{\alpha}\partial_{\beta}a_{0\gamma}}-\frac{c_{P}^{2}+c_{S}^{2}}{c_{S}^{2}}\partial_{\alpha}(c_{S}^{2})g^{\beta\gamma}\partial_{\beta}a_{0\gamma}+(c_{P}^{2}-c_{S}^{2})\partial_{\alpha}g^{\beta\gamma}\partial_{\beta}a_{0\gamma}$
$+2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\partial_{\alpha}a_{0\gamma}+2\partial_{\alpha}g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$+2g^{\beta\gamma}c_{S}^{2}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$+(c_{P}^{2}+3c_{S}^{2})c_{S}^{-1}g^{\beta\gamma}\partial_{\beta}c_{S}\partial_{\alpha}a_{0\gamma}-{(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\Gamma^{\delta}_{\beta\gamma}\partial_{\alpha}a_{0\delta}}$
$+2g^{\beta\gamma}\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}-2\frac{c_{S}^{2}\partial_{\alpha}c_{P}^{2}+c_{P}^{2}\partial_{\alpha}c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$-2c_{P}^{2}c_{S}^{-1}\partial_{\alpha}c_{S}\frac{1}{c_{P}^{2}-c_{S}^{2}}\left((c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\partial_{\beta}a_{0\gamma}+2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}\right)$
$+\partial_{\alpha}(c_{P}^{2}-2c_{S}^{2})\frac{1}{c_{P}^{2}-c_{S}^{2}}\left({(c_{P}^{2}-c_{S}^{2})g^{\beta\gamma}\partial_{\beta}a_{0\gamma}}+2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}\right)$
$+\partial_{\alpha}\log\sqrt{\rho}\frac{2(c_{P}^{2}-2c_{S}^{2})}{c_{P}^{2}-c_{S}^{2}}\left({\frac{\lambda+\mu}{\rho}g^{\beta\gamma}\partial_{\beta}a_{0\gamma}}+g^{\gamma\beta}\frac{\partial_{\beta}\mu}{\rho}a_{0\gamma}+\frac{\lambda+3\mu}{\rho}c_{S}^{-1}g^{\beta\gamma}\partial_{\beta}c_{S}a_{0\gamma}-{\frac{\lambda+\mu}{\rho}g^{\beta\gamma}\Gamma^{\delta}_{\beta\gamma}a_{0\delta}}\right)$
$-{\rho^{-1}\partial_{\alpha}\lambda g^{\delta\varepsilon}\partial_{\varepsilon}a_{0\delta}}-{(c_{P}^{2}-2c_{S}^{2})g^{\delta\varepsilon}\partial_{\alpha}\partial_{\varepsilon}a_{0\delta}}-(c_{P}^{2}-2c_{S}^{2})\partial_{\varepsilon}a_{0\delta}\partial_{\alpha}g^{\delta\varepsilon}$
$-\rho^{-1}\partial_{\alpha}\lambda c_{S}^{-1}g^{\delta\beta}\partial_{\beta}c_{S}a_{0\delta}-(c_{P}^{2}-2c_{S}^{2})c_{S}^{-1}g^{\delta\beta}\partial_{\beta}c_{S}\partial_{\alpha}a_{0\delta}+{\rho^{-1}\partial_{\alpha}\lambda g^{\delta\varepsilon}\Gamma^{\eta}_{\varepsilon\delta}a_{0\eta}}+{(c_{P}^{2}-2c_{S}^{2})g^{\delta\varepsilon}\Gamma^{\eta}_{\varepsilon\delta}\partial_{\alpha}a_{0\eta}}$
$-c_{S}^{-2}\rho^{-1}\partial_{3}\mu\partial_{3}a_{0\alpha}-\partial^{2}_{33}a_{0\alpha}-\rho^{-1}g^{\beta\gamma}\partial_{\beta}\mu\partial_{\gamma}a_{0\alpha}-c_{S}^{2}g^{\beta\gamma}\partial_{\beta}\partial_{\gamma}a_{0\alpha}-g^{\beta\gamma}\rho^{-1}\partial_{\gamma}\mu\partial_{\alpha}a_{0\beta}-{\mu g^{\beta\gamma}\partial_{\alpha}\partial_{\gamma}a_{0\beta}}$
$+2\Gamma^{\beta}_{\alpha 3}\partial_{3}a_{0\beta}+2g^{\delta\gamma}c_{S}^{2}\Gamma^{\beta}_{\alpha\gamma}\partial_{\delta}a_{0\beta}+\partial_{3}\rho g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}a_{0\beta}+2c_{S}^{2}g^{\delta\gamma}\partial_{\delta}\rho\Gamma^{\beta}_{\alpha\gamma}a_{0\beta}$
$+c_{S}^{-1}\partial_{3}c_{S}\partial_{3}a_{0\alpha}-\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}\partial_{3}a_{0\alpha}$
$+g^{\gamma\delta}\Gamma^{\beta}_{\alpha\gamma}c_{S}^{2}\partial_{\delta}a_{0\beta}+g^{\beta\gamma}\Gamma^{\delta}_{\alpha\gamma}c_{S}^{2}\partial_{\delta}a_{0\beta}$
$+c_{S}^{2}(-c_{S}^{-1}g^{\beta\gamma}\partial_{\gamma}c_{S})(\partial_{\alpha}a_{0\beta}+\partial_{\beta}a_{0\alpha})+c_{S}^{2}g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta}({\partial_{\alpha}a_{0\beta}}+\partial_{\beta}a_{0\alpha})$
$+\frac{1}{\sqrt{\rho}}\times\text{terms that do not depend on $\rho$}.$
Here we have also used the identity (8).
Tedious calculations give
$H=$ $-\partial^{2}_{33}a_{0\alpha}-c_{S}^{-2}\rho^{-1}\partial_{3}\mu\partial_{3}a_{0\alpha}+2\Gamma^{\beta}_{\alpha 3}\partial_{3}a_{0\beta}$
$-g^{\beta\gamma}\rho^{-1}\partial_{\beta}\mu\partial_{\gamma}a_{0\alpha}-c_{S}^{2}g^{\beta\gamma}\partial_{\beta}\partial_{\gamma}a_{0\alpha}-\rho^{-1}g^{\beta\gamma}\partial_{\gamma}\mu\partial_{\alpha}a_{0\beta}$
$+2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\partial_{\alpha}a_{0\gamma}+2g^{\beta\gamma}c_{S}^{2}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$+2(c_{P}^{2}-2c_{S}^{2})\partial_{\alpha}\log\sqrt{\rho}\frac{1}{c_{P}^{2}-c_{S}^{2}}g^{\gamma\beta}\frac{\partial_{\beta}\mu}{\rho}a_{0\gamma}$
$-\frac{2c_{P}^{2}+c_{S}^{2}}{c_{S}^{2}}\partial_{\alpha}(c_{S}^{2})g^{\beta\gamma}\partial_{\beta}a_{0\gamma}-\partial_{\alpha}(c_{P}^{2}-2c_{S}^{2})g^{\delta\varepsilon}\partial_{\varepsilon}a_{0\delta}$
$+(c_{P}^{2}-c_{S}^{2})\partial_{\alpha}g^{\beta\gamma}\partial_{\beta}a_{0\gamma}+2\partial_{\alpha}g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}-(c_{P}^{2}-2c_{S}^{2})\partial_{\varepsilon}a_{0\delta}\partial_{\alpha}g^{\delta\varepsilon}$
$+3g^{\delta\gamma}c_{S}^{2}\Gamma^{\beta}_{\alpha\gamma}\partial_{\delta}a_{0\beta}+g^{\beta\gamma}\Gamma^{\delta}_{\alpha\gamma}c_{S}^{2}\partial_{\delta}a_{0\beta}$
$+2\partial_{3}\log\sqrt{\rho}g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}a_{0\beta}+4c_{S}^{2}g^{\delta\gamma}\partial_{\delta}\log\sqrt{\rho}\Gamma^{\beta}_{\alpha\gamma}a_{0\beta}$
$+(-2\partial_{\alpha}c_{S}^{2}-\frac{8c_{S}^{2}\partial c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}})g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$+2(c_{P}^{2}-2c_{S}^{2})\partial_{\alpha}\log\sqrt{\rho}\frac{1}{c_{P}^{2}-c_{S}^{2}}\frac{\lambda+3\mu}{\rho}c_{S}^{-1}g^{\beta\gamma}\partial_{\beta}c_{S}a_{0\gamma}-\rho^{-1}\partial_{\alpha}\lambda c_{S}^{-1}g^{\delta\beta}\partial_{\beta}c_{S}a_{0\delta}$
$-c_{S}\partial_{\gamma}c_{S}g^{\beta\gamma}\partial_{\beta}a_{0\alpha}+\mu g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta}\partial_{\beta}a_{0\alpha}$
$+c_{S}^{-3}\partial_{3}c_{S}\mu\partial_{3}a_{0\alpha}-\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}\partial_{3}a_{0\alpha}$
$+4c_{S}g^{\beta\gamma}\partial_{\beta}c_{S}\partial_{\alpha}a_{0\gamma}$
$+\frac{1}{\sqrt{\rho}}\times\text{terms that do not depend on $\rho$}.$
Recalling $a_{0\alpha}=A_{S}c_{S}\eta_{\alpha}$, we have
$H=$ $-\partial^{2}_{33}A_{S}c_{S}\eta_{\alpha}{-2\partial_{3}A_{S}\partial_{3}(c_{S}\eta_{\alpha}})-c_{S}^{-1}\partial_{3}c_{S}^{2}\partial_{3}A_{S}\eta_{\alpha}-2\partial_{3}\log\sqrt{\rho}\partial_{3}A_{S}c_{S}\eta_{\alpha}$
${-2\partial_{3}(c_{S}\log\sqrt{\rho}A_{S}\partial_{3}\eta_{\alpha}})+g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}\partial_{3}A_{S}c_{S}\eta_{\beta}$
$-g^{\beta\gamma}\partial_{\beta}c_{S}^{2}\partial_{\gamma}A_{S}c_{S}\eta_{\alpha}-2g^{\beta\gamma}c_{S}^{3}\partial_{\beta}\log\sqrt{\rho}\partial_{\gamma}A_{S}\eta_{\alpha}-{2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}A_{S}\partial_{\gamma}(c_{S}\eta_{\alpha}})$
$-\mu g^{\beta\gamma}\partial_{\beta}\partial_{\gamma}A_{S}c_{S}\eta_{\alpha}-{2\mu g^{\beta\gamma}\partial_{\beta}A_{S}\partial_{\gamma}(c_{S}\eta_{\alpha}})$
$-g^{\beta\gamma}\partial_{\gamma}c_{S}^{2}\partial_{\alpha}A_{S}c_{S}\eta_{\beta}-{2g^{\beta\gamma}c_{S}^{2}\partial_{\gamma}\log\sqrt{\rho}\partial_{\alpha}A_{S}c_{S}\eta_{\beta}}-{2g^{\beta\gamma}c_{S}^{2}\partial_{\gamma}\log\sqrt{\rho}A_{S}\partial_{\alpha}(c_{S}\eta_{\beta}})$
$+{2g^{\beta\gamma}c_{S}^{3}\partial_{\beta}\log\sqrt{\rho}\partial_{\alpha}A_{S}\eta_{\gamma}}+{2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}A_{S}\partial_{\alpha}(c_{S}\eta_{\gamma}})+2g^{\beta\gamma}c_{S}^{3}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}a_{0\gamma}$
$+2\partial_{\alpha}\log\sqrt{\rho}\frac{c_{P}^{2}-2c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}g^{\gamma\beta}\partial_{\beta}c_{S}^{2}A_{S}c_{S}\eta_{\gamma}+4\partial_{\alpha}\log\sqrt{\rho}\frac{c_{P}^{2}-2c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}g^{\gamma\beta}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}A_{S}c_{S}\eta_{\gamma}$
$-\frac{2c_{P}^{2}+c_{S}^{2}}{c_{S}^{2}}\partial_{\alpha}(c_{S}^{2})g^{\beta\gamma}\partial_{\beta}A_{S}c_{S}\eta_{\gamma}-\partial_{\alpha}(c_{P}^{2}-2c_{S}^{2})g^{\delta\varepsilon}\partial_{\varepsilon}A_{S}c_{S}\eta_{\delta}$
$+(c_{P}^{2}-c_{S}^{2})\partial_{\alpha}g^{\beta\gamma}\partial_{\beta}A_{S}c_{S}\eta_{\gamma}+2\partial_{\alpha}g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}c_{S}A_{S}\eta_{\gamma}$
$-\lambda\partial_{\varepsilon}A_{S}c_{S}\eta_{\delta}\partial_{\alpha}g^{\delta\varepsilon}+3g^{\delta\gamma}\mu\Gamma^{\beta}_{\alpha\gamma}\partial_{\delta}A_{S}c_{S}\eta_{\beta}+g^{\beta\gamma}\Gamma^{\delta}_{\alpha\gamma}\mu\partial_{\delta}A_{S}c_{S}\eta_{\beta}$
$+(-2\partial_{\alpha}c_{S}^{2}-\frac{8c_{S}^{2}\partial c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}})g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}A_{S}c_{S}\eta_{\gamma}$
$+2(c_{P}^{2}-2c_{S}^{2})\partial_{\alpha}\log\sqrt{\rho}\frac{1}{c_{P}^{2}-c_{S}^{2}}\frac{\lambda+3\mu}{\rho}c_{S}^{-1}g^{\beta\gamma}\partial_{\beta}c_{S}^{2}A_{S}\eta_{\gamma}-2\partial_{\alpha}\log\sqrt{\rho}c_{S}g^{\delta\beta}\partial_{\beta}c^{2}_{S}A_{S}\eta_{\delta}$
$-c_{S}\partial_{\gamma}c_{S}g^{\beta\gamma}\partial_{\beta}A_{S}c_{S}\eta_{\alpha}+c_{S}^{2}g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta}\partial_{\beta}A_{S}c_{S}\eta_{\alpha}$
$+c_{S}^{-3}\partial_{3}c_{S}\mu c_{S}\partial_{3}A_{S}\eta_{\alpha}-\frac{1}{2}g^{\beta\gamma}c_{S}^{-2}\frac{\partial g_{\beta\gamma}}{\partial\tau}\mu c_{S}\partial_{3}A_{S}\eta_{\alpha}$
$+2\partial_{3}\log\sqrt{\rho}g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}c_{S}A_{S}\eta_{\beta}+4c_{S}^{2}g^{\delta\gamma}\partial_{\delta}\log\sqrt{\rho}\Gamma^{\beta}_{\alpha\gamma}c_{S}A_{S}\eta_{\beta}$
$+4c_{S}g^{\beta\gamma}\partial_{\beta}c_{S}\partial_{\alpha}c_{S}A_{S}\eta_{\gamma}$
$+\frac{1}{\sqrt{\rho}}\times\text{terms that do not depend on $\rho$}.$
Upon using
$\partial A_{S}=-J^{-1/2}c_{S}^{-1/2}\rho^{-1/2}\partial\log\sqrt{\rho}-\frac{1}{2}c_{S}^{-3/2}\partial_{3}c_{S}J^{-1/2}\rho^{-1/2}-\frac{1}{2}J^{-3/2}\partial Jc_{S}^{-1/2}\rho^{-1/2},$
$\begin{split}\partial^{2}_{33}A_{S}=&-J^{-1/2}c_{S}^{-1/2}\rho^{-1/2}\partial^{2}_{3}\log\sqrt{\rho}+J^{-1/2}c_{S}^{-1/2}\rho^{-1/2}(\partial_{3}\log\sqrt{\rho})^{2}\\ &+c_{S}^{-3/2}\partial_{3}c_{S}J^{-1/2}\rho^{-1/2}\partial_{3}\log\sqrt{\rho}+J^{-3/2}\partial_{3}Jc_{S}^{-1/2}\rho^{-1/2}\partial_{3}\log\sqrt{\rho}\\ &+\frac{1}{\sqrt{\rho}}\times\text{terms that do not depend on $\rho$},\end{split}$
and
$\begin{split}g^{\beta\gamma}\partial_{\beta}\partial_{\gamma}A_{S}=&J^{-1/2}c_{S}^{-1/2}g^{\beta\gamma}\partial_{\beta}\partial_{\gamma}\rho^{-1/2}+2g^{\beta\gamma}\partial_{\beta}J^{-1/2}c_{S}^{-1/2}\partial_{\gamma}\rho^{-1/2}+2J^{-1/2}g^{\beta\gamma}\partial_{\beta}c_{S}^{-1/2}\partial_{\gamma}\rho^{-1/2}\\ =&-J^{-1/2}c_{S}^{-1/2}\rho^{-1/2}g^{\beta\gamma}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}+J^{-1/2}c_{S}^{-1/2}\rho^{-1/2}g^{\beta\gamma}(\partial_{\alpha}\log\sqrt{\rho})(\partial_{\beta}\log\sqrt{\rho})\\ &+g^{\beta\gamma}J^{-3/2}\partial_{\beta}Jc_{P}^{-1/2}\rho^{-1/2}\partial_{\gamma}\log\sqrt{\rho}+J^{-1/2}\rho^{-1/2}g^{\beta\gamma}c_{S}^{-3/2}\partial_{\beta}c_{S}\partial_{\gamma}\log\sqrt{\rho}\\ &+\frac{1}{\sqrt{\rho}}\times\text{terms that do not depend on $\rho$},\end{split}$
we have
$\rho H=\rho^{1/2}c_{S}^{1/2}J^{-1/2}H^{\prime},$
where
$H^{\prime}=$ $\partial^{2}_{3}\log\sqrt{\rho}\eta_{\alpha}-(\partial_{3}\log\sqrt{\rho})^{2}\eta_{\alpha}-c_{S}^{-1}\partial_{3}c_{S}\partial_{3}\log\sqrt{\rho}\eta_{\alpha}-{J^{-1}\partial_{3}J\partial_{3}\log\sqrt{\rho}\eta_{\alpha}}+c_{S}^{-2}\partial_{3}c_{S}^{2}\partial_{3}\log\sqrt{\rho}\eta_{\alpha}$
$-2\partial_{3}\log\sqrt{\rho}\left(-\partial_{3}\log\sqrt{\rho}-\frac{1}{2}c_{S}^{-1}\partial_{3}c_{S}-{\frac{1}{2}J^{-1}\partial_{3}J}\right)\eta_{\alpha}+g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}\partial_{3}\log\sqrt{\rho}\eta_{\beta}$
$+g^{\beta\gamma}\partial_{\beta}c_{S}^{2}\partial_{\gamma}\log\sqrt{\rho}\eta_{\alpha}-2g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\left(-\partial_{\gamma}\log\sqrt{\rho}-\frac{1}{2}c_{S}^{-1}\partial_{\gamma}c_{S}-{\frac{1}{2}J^{-1}\partial_{\gamma}J}\right)\eta_{\alpha}$
$+c_{S}^{2}g^{\beta\gamma}\partial_{\gamma}\partial_{\beta}\log\sqrt{\rho}\eta_{\alpha}-c_{S}^{2}g^{\beta\gamma}(\partial_{\gamma}\log\sqrt{\rho})(\partial_{\beta}\log\sqrt{\rho})\eta_{\alpha}-{c_{S}^{2}g^{\beta\gamma}J^{-1}\partial_{\beta}J\partial_{\gamma}\log\sqrt{\rho}\eta_{\alpha}}$
$-c_{S}^{2}g^{\beta\gamma}c_{S}^{-1}\partial_{\beta}c_{S}\partial_{\gamma}\log\sqrt{\rho}\eta_{\alpha}$
$+g^{\beta\gamma}\partial_{\gamma}c_{S}^{2}\partial_{\alpha}\log\sqrt{\rho}\eta_{\beta}+2g^{\beta\gamma}c_{S}^{2}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}$
$+2(c_{P}^{2}-2c_{S}^{2})\partial_{\alpha}\log\sqrt{\rho}\frac{1}{c_{P}^{2}-c_{S}^{2}}g^{\gamma\beta}\partial_{\beta}c_{S}^{2}\eta_{\gamma}+4(c_{P}^{2}-2c_{S}^{2})\partial_{\alpha}\log\sqrt{\rho}\frac{1}{c_{P}^{2}-c_{S}^{2}}g^{\gamma\beta}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}$
$+\frac{2c_{P}^{2}+c_{S}^{2}}{c_{S}^{2}}\partial_{\alpha}(c_{S}^{2})g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}+\partial_{\alpha}(c_{P}^{2}-2c_{S}^{2})g^{\delta\varepsilon}\partial_{\varepsilon}\log\sqrt{\rho}\eta_{\delta}$
$-c_{S}^{2}\partial_{\alpha}g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}+2\partial_{\alpha}g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}+g^{\delta\gamma}c_{S}^{2}\Gamma^{\beta}_{\alpha\gamma}\partial_{\delta}\log\sqrt{\rho}\eta_{\beta}-g^{\beta\gamma}\Gamma^{\delta}_{\alpha\gamma}c_{S}^{2}\partial_{\delta}\log\sqrt{\rho}\eta_{\beta}$
$+(-2\partial_{\alpha}c_{S}^{2}-\frac{8c_{S}^{2}\partial_{\alpha}c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}})g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}$
$+2(c_{P}^{2}-2c_{S}^{2})\partial_{\alpha}\log\sqrt{\rho}\frac{1}{c_{P}^{2}-c_{S}^{2}}(c_{P}^{2}+c_{S}^{2})c_{S}^{-1}g^{\beta\gamma}\partial_{\beta}c_{S}\eta_{\gamma}-2\partial_{\alpha}\log\sqrt{\rho}c_{S}g^{\delta\beta}\partial_{\beta}c_{S}\eta_{\delta}$
$+c_{S}\partial_{\gamma}c_{S}g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}\eta_{\alpha}-c_{S}^{2}g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta}\partial_{\beta}\log\sqrt{\rho}\eta_{\alpha}$
$-c_{S}^{-3}\partial_{3}c_{S}c_{S}^{2}\partial_{3}\log\sqrt{\rho}\eta_{\alpha}+\frac{1}{2}g^{\beta\gamma}c_{S}^{-2}\frac{\partial g_{\beta\gamma}}{\partial\tau}c_{S}^{2}\partial_{3}\log\sqrt{\rho}\eta_{\alpha}-4c_{S}^{2}g^{\beta\gamma}\partial_{\beta}c_{S}\partial_{\alpha}\log\sqrt{\rho}\eta_{\gamma}$
+terms that do not depend on $+\text{terms that do not depend on $\rho$}.$
Note that
$\begin{split}&\partial_{\alpha}g^{\beta\gamma}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}+g^{\delta\gamma}c_{S}^{2}\Gamma^{\beta}_{\alpha\gamma}\partial_{\delta}\log\sqrt{\rho}\eta_{\beta}-g^{\beta\gamma}\Gamma^{\delta}_{\alpha\gamma}c_{S}^{2}\partial_{\delta}\log\sqrt{\rho}\eta_{\beta}\\ =&c_{S}^{2}\left(-g^{\beta\delta}\partial_{\alpha}g_{\delta\varepsilon}g^{\varepsilon\gamma}+\frac{1}{2}g^{\beta\delta}g^{\gamma\varepsilon}(\partial_{\alpha}g_{\delta\varepsilon}+\partial_{\delta}g_{\alpha\varepsilon}-\partial_{\varepsilon}g_{\alpha\delta})-\frac{1}{2}g^{\gamma\delta}g^{\beta\varepsilon}(\partial_{\alpha}g_{\delta\varepsilon}+\partial_{\delta}g_{\alpha\varepsilon}-\partial_{\varepsilon}g_{\alpha\delta})\right)\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}\\ =&-c_{S}^{2}g^{\gamma\delta}g^{\beta\varepsilon}(\partial_{\alpha}g_{\delta\varepsilon}+\partial_{\delta}g_{\alpha\varepsilon}-\partial_{\varepsilon}g_{\alpha\delta})\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}\\ =&-2c_{S}^{2}g^{\gamma\delta}\Gamma^{\beta}_{\alpha\delta}\partial_{\beta}\log\sqrt{\rho}\eta_{\gamma}.\end{split}$
Notice also that
$c_{S}^{-2}|\eta|^{2}=1,$
that is $g_{\alpha\beta}\eta^{\alpha}\eta^{\beta}=g^{\alpha\beta}\eta_{\alpha}\eta_{\beta}=c_{S}^{2}$. Using also (11) and (12), we then have
$c_{S}^{-2}H^{\prime}_{\alpha}\eta^{\alpha}$
$=$ $\partial^{2}_{3}\log\sqrt{\rho}+(\partial_{3}\log\sqrt{\rho})^{2}+c_{S}^{2}g^{\beta\gamma}(\partial_{\gamma}\log\sqrt{\rho})(\partial_{\beta}\log\sqrt{\rho})+c_{S}^{2}g^{\beta\gamma}\partial_{\gamma}\partial_{\beta}\log\sqrt{\rho}+$
$+2\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}\eta^{\beta}\eta^{\alpha}+4(c_{P}^{2}-2c_{S}^{2})\frac{1}{c_{P}^{2}-c_{S}^{2}}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}\log\sqrt{\rho}\eta^{\beta}\eta^{\alpha}-c_{S}^{2}g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta}\partial_{\beta}\log\sqrt{\rho}$
$+\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}\partial_{3}\log\sqrt{\rho}+c_{S}^{-2}\frac{\partial g_{\alpha\gamma}}{\partial\tau}\partial_{3}\log\sqrt{\rho}\eta^{\gamma}\eta^{\alpha}-2\Gamma^{\beta}_{\alpha\delta}\partial_{\beta}\log\sqrt{\rho}\eta^{\delta}\eta^{\alpha}$
$+c_{S}^{-1}\partial_{3}c_{S}\partial_{3}\log\sqrt{\rho}+3c_{S}g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}\partial_{\gamma}c_{S}$
$+3\frac{c_{P}^{2}-c_{S}^{2}}{c_{S}^{4}}\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}+c_{S}^{-2}\partial_{\alpha}c_{P}^{2}\partial_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}-\frac{12}{c_{P}^{2}-c_{S}^{2}}\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}$
+terms that do not depend on $+\text{terms that do not depend on $\rho$}$
$=$ $\partial^{2}_{3}\log\sqrt{\rho}-c_{S}^{-1}\partial_{3}c_{S}\partial_{3}\log\sqrt{\rho}+c_{S}g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}\partial_{\gamma}c_{S}+c_{S}^{2}g^{\beta\gamma}\partial_{\gamma}\partial_{\beta}\log\sqrt{\rho}$
$+\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}\partial_{3}\log\sqrt{\rho}-c_{S}^{2}g^{\eta\varepsilon}\Gamma_{\eta\varepsilon}^{\beta}\partial_{\beta}\log\sqrt{\rho}$
$+2\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}\eta^{\beta}\eta^{\alpha}+c_{S}^{-2}\frac{\partial g_{\alpha\gamma}}{\partial\tau}\partial_{3}\log\sqrt{\rho}\eta^{\gamma}\eta^{\alpha}-2\Gamma^{\beta}_{\alpha\delta}\partial_{\beta}\log\sqrt{\rho}\eta^{\delta}\eta^{\alpha}$
$+c_{S}^{2}g^{\beta\gamma}(\partial_{\gamma}\log\sqrt{\rho})(\partial_{\beta}\log\sqrt{\rho})+(\partial_{3}\log\sqrt{\rho})^{2}$
$+4(c_{P}^{2}-2c_{S}^{2})\frac{1}{c_{P}^{2}-c_{S}^{2}}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}\log\sqrt{\rho}\eta^{\beta}\eta^{\alpha}$
$+c_{S}^{-2}\partial_{3}c_{S}^{2}\partial_{3}\log\sqrt{\rho}+g^{\beta\gamma}\partial_{\beta}\log\sqrt{\rho}\partial_{\gamma}c_{S}^{2}$
$+3\frac{c_{P}^{2}-c_{S}^{2}}{c_{S}^{4}}\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}+c_{S}^{-2}\partial_{\alpha}c_{P}^{2}\partial_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}-\frac{12}{c_{P}^{2}-c_{S}^{2}}\partial_{\alpha}c_{S}^{2}\partial_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}$
+terms that do not depend on $+\text{terms that do not depend on $\rho$}$
$=$ $c_{S}^{2}\Delta\log\sqrt{\rho}+2\nabla_{\alpha}\nabla_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}+c_{S}^{2}|\nabla\log\sqrt{\rho}|^{2}+4\frac{c_{P}^{2}-2c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}\nabla_{\alpha}\log\sqrt{\rho}\nabla_{\beta}\log\sqrt{\rho}\eta^{\alpha}\eta^{\beta}$
$+\nabla\log\sqrt{\rho}\cdot\nabla c_{S}^{2}+\nabla_{\beta}\log\sqrt{\rho}\left(2\frac{c_{P}^{2}-c_{S}^{2}}{c_{S}^{4}}\nabla_{\alpha}c_{S}^{2}+c_{S}^{-2}\nabla_{\alpha}c_{P}^{2}-\frac{12}{c_{P}^{2}-c_{S}^{2}}\nabla_{\alpha}c_{S}^{2}\right)\eta^{\alpha}\eta^{\beta}$
+terms that do not depend on $+\text{terms that do not depend on $\rho$}$
$=$ $:N_{\alpha\beta}\eta^{\alpha}\eta^{\beta}+C,$
where $C$ does not depend on $\rho$. Note that the second order tensor $N$ can be written in a coordinate-free way
$\begin{split}N=&c_{S}^{2}\Delta\log\sqrt{\rho}g_{S}+2\nabla^{2}\log\sqrt{\rho}+c_{S}^{2}|\nabla\log\sqrt{\rho}|^{2}g_{S}+4\frac{c_{P}^{2}-2c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}\nabla\log\sqrt{\rho}\otimes\nabla\log\sqrt{\rho}\\ &+(\nabla\log\sqrt{\rho}\cdot\nabla c_{S}^{2})g_{S}+\nabla\log\sqrt{\rho}\otimes\left(2\frac{c_{P}^{2}-c_{S}^{2}}{c_{S}^{4}}\nabla c_{S}^{2}+c_{S}^{-2}\nabla c_{P}^{2}-\frac{12}{c_{P}^{2}-c_{S}^{2}}\nabla c_{S}^{2}\right).\end{split}$
Multiplying both sides of (14) by $c_{S}^{2}\eta^{\alpha}$, we get
$\begin{split}&\mathrm{i}c_{S}^{-2}\left(2\mu\partial_{3}a_{1\alpha}+(\partial_{3}\mu)a_{1\alpha}-\mu g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}a_{1\beta}+\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}c_{S}^{-2}\mu a_{1\alpha}-c_{S}^{-1}\partial_{3}c_{S}\mu a_{1\alpha}\right)c_{S}^{2}\eta^{\alpha}\\ =&(\langle N,\eta\otimes\eta\rangle_{g_{S}}+C)J^{-1/2}c_{S}^{1/2}\rho^{1/2}.\end{split}$
Assume that $\gamma$ is an arbitrary geodesic in $(\Omega,g_{S})$ connecting two points $a\in\partial\Omega$ to $b\in\partial\Omega$, so the length of $\gamma$ is smaller than $T$. Denote $D_{\tau}$ to be the covariant derivative along the geodesic. We consider the above equation along $\gamma$. If we denote
$a_{1\alpha}=J^{-1/2}c_{S}^{1/2}\rho^{-1/2}\xi_{\alpha},$
with a vector field $\xi=(\xi_{1},\xi_{2},0)$ along $\gamma$, the above equation can be written as
$D_{\tau}\langle\xi,\eta\rangle_{g_{S}}=\langle D_{\tau}\xi,\eta\rangle_{g_{S}}=-\frac{\mathrm{i}}{2}\langle N,\eta\otimes\eta\rangle_{g_{S}}+C$
along $\gamma$.
Then
$\langle\xi,\eta\rangle_{g_{S}}\Big{|}_{a}^{b}=c_{S}^{2}(\xi\cdot\eta)\Big{|}_{a}^{b}=-\frac{\mathrm{i}}{2}\int_{\gamma}(\langle N,\eta\otimes\eta\rangle_{g_{S}}+C)\mathrm{d}s.$
So we can recover the transverse ray transform of $N$
$\int_{\gamma}\langle N(\gamma(s)),\eta\otimes\eta\rangle_{g_{S}}\mathrm{d}s$
from boundary value of $\mathbf{a}_{1}$ once the boundary value of $\rho$ is determined. Here $\eta$ could be an arbitrary parallel vector field along $\gamma$ and orthogonal to $\gamma$ with respect the metric $g_{S}$ (also with respect to the Euclidean metric $g_{E}$). Thus the transverse ray transform of $N$ can be recovered from the Dirichlet-to-Neumann map $\Lambda_{\rho}$. We note here that clearly the transverse ray transform of $N$ is equal to that of the symmetrization of $N$.
Next we work under the Cartesian coordinates. Under the strictly convex foliation condition, by the invertibility of the transverse ray transform [33], we can recover the symmetrization of $N$,
$\begin{split}\mathrm{Sym}(N)=&\Delta\log\sqrt{\rho}I+2\nabla^{2}\log\sqrt{\rho}+|\nabla\log\sqrt{\rho}|^{2}I+4\frac{c_{P}^{2}-2c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}\nabla\log\sqrt{\rho}\otimes\nabla\log\sqrt{\rho}\\ &+c_{S}^{-2}(\nabla\log\sqrt{\rho}\cdot\nabla c_{S}^{2})I+\nabla\log\sqrt{\rho}\otimes^{s}\left(2\frac{c_{P}^{2}-c_{S}^{2}}{c_{S}^{4}}\nabla c_{S}^{2}+c_{S}^{-2}\nabla c_{P}^{2}-\frac{12}{c_{P}^{2}-c_{S}^{2}}\nabla c_{S}^{2}\right).\end{split}$
Here $I$ is the identity matrix (the Euclidean metric under Cartesian coordinates) and $\otimes^{s}$ denotes the symmetrized product
$u\otimes^{s}v=\frac{1}{2}(u\otimes v+v\otimes u).$
We also refer the readers to [21] for the invertibility of the transverse ray transforms under certain curvature conditions. In conclusion, if $\Lambda_{\rho_{1}}=\Lambda_{\rho_{2}}$, we have $\mathrm{Sym}(N_{1})=\mathrm{Sym}(N_{2})$, which can be written as a second order differential equation for $\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}}$,
$\begin{split}&\Delta(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}})I+2\nabla^{2}(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}})+(|\nabla\log\sqrt{\rho_{1}}|^{2}-|\nabla\log\sqrt{\rho_{2}}|^{2})I\\ &+4\frac{c_{P}^{2}-2c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}\left(\nabla\log\sqrt{\rho_{1}}\otimes\nabla\log\sqrt{\rho_{1}}-\nabla\log\sqrt{\rho_{2}}\otimes\nabla\log\sqrt{\rho_{2}}\right)+c_{S}^{-2}(\nabla(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}})\cdot\nabla c_{S}^{2})I\\ &+\left(2\frac{c_{P}^{2}-c_{S}^{2}}{c_{S}^{4}}\nabla c_{S}^{2}+c_{S}^{-2}\nabla c_{P}^{2}-\frac{12}{c_{P}^{2}-c_{S}^{2}}\nabla c_{S}^{2}\right)\otimes^{s}(\nabla\log\sqrt{\rho_{1}}-\nabla\log\sqrt{\rho_{2}})\\ =&0\quad\text{in }\Omega.\end{split}$
Applying divergence twice to the above equation, we obtain a fourth order linear differential equation of the form
$3\Delta^{2}(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}})+\sum_{|\alpha|\leq 3}F_{\alpha}(\rho_{1},\rho_{2})D^{\alpha}(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}})=0,\quad\text{in }\Omega.$
By the Unique Continuation Principle (cf. [17]) and the boundary determination results $\partial_{\nu}^{k}\rho_{1}=\partial_{\nu}^{k}\rho_{2}$, $k=0,1,2,3$ on $\partial\Omega$ (cf. [18]), we can conclude that
$\rho_{1}=\rho_{2}\quad\text{in }\overline{\Omega}.$
This completes the proof of the main theorem.